### v1

# Varco Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models

Abstract
The rapid advancement of Large Language Models (LLMs) necessitates robust evaluation methodologies. Current benchmarking approaches often rely on comparing model outputs against predefined prompts and reference outputs. Relying on predefined reference outputs hinders flexible adaptation of benchmarks to the rapidly evolving capabilities of LLMs. This limitation necessitates periodic efforts to prepare new benchmarks. To keep pace with rapidly evolving LLM capabilities, we propose a more flexible benchmarking approach. Our method, Varco Arena, provides reference-free benchmarking of LLMs in tournament style. Varco Arena directly compares LLM outputs across a diverse set of prompts, determining model rankings through a single-elimination tournament structure. This direct pairwise comparison offers two key advantages: (1) Direct comparison, unmediated by reference text, more effectively orders competing LLMs, resulting in more reliable rankings, and (2) reference-free approach to benchmarking adds flexibility in updating benchmark prompts by eliminating the need for quality references. Our empirical results, supported by simulation experiments, demonstrate that the Varco Arena tournament approach aligns better with the current Elo model for benchmarking LLMs. The alignment is measured in terms of Spearman correlation, showing improvement over current practice of benchmarking that use reference outputs as comparison anchors.
1 Introduction
The versatility of Large Language Models (LLMs) stems from their generative capacity to address a wide array of tasks. The multi-faceted capability of LLMs enables flexible applications across numerous user scenarios, exemplified by generalist assistants (Ouyang et al. 2022; Köpf et al. 2024; Roziere et al. 2023). As versatility emerges as a core attribute of LLMs, the challenge of accurately gauging their skill becomes increasingly significant. In response to this challenge, numerous benchmarks evaluating LLM capability have emerged (Hendrycks et al. 2020; Srivastava et al. 2023; Zhong et al. 2024). Many LLM benchmarks employ formats amenable to automated scoring. Examples include benchmarks for arithmetic problems (Gao et al. 2022; Patel, Bhattamishra, and Goyal 2021), multiple-choice questions (Lin, Hilton, and Evans 2022), and code execution (Austin et al. 2021; Chen et al. 2021). While these benchmarks are valuable, they primarily assess problem-solving abilities. The need for LLM benchmark extends beyond this, as the primary value of LLMs lies in their generative capabilities. Researchers have primarily relied on pariwise comparison to evaluate and rank the generative capabilities of LLMs. A representative example of this approach is Chatbot Arena (Chiang et al. 2024), which computes Elo ratings based on a massive number of human-annotated pairwise comparisons. While Chatbot Arena benefited from the reliability of open-ended prompts from users, acquiring such a large number of annotations for benchmarking LLMs is time-consuming and resource-intensive. To address these challenges, several benchmark datasets for generation tasks have been developed. These leverage reference responses for ranking LLMs using automated LLM judges (Zheng et al. 2024) includes AlpacaEval (Li et al. 2023), Arena-Hard-Auto (Li et al. 2024), and MTBench (Zheng et al. 2023). Relying on reference outputs offers two key advantages: (1) The number of comparisons required for ranking multiple LLMs reduces linear to the number of LLMs. (2) Reference texts establish a quality bar for evaluating responses.
However, we propose that directly comparing the responses of LLMs, rather than relying on reference outputs, is a more effective way to achieve accurate LLM rankings. We suggest that combined with tournament-styled match making, direct comparison of outputs can achieve a better correlation to human preference within the same number of comparisons. Additionally, reference-free benchmarking mitigates potential biases in open-ended generation tasks and provides more flexibility compared to static benchmark datasets.
Figure 1: Varco Arena directly compares LLM output pairs in single-elimination tournament structure rather than comparing LLM outputs to reference text. In terms of deciding whether a certain LLM is better or worse compared to the other one, we suggest direct comparison is more intuitive and results in better separability.
We devised Varco Arena, which directly compares LLM outputs without the need for reference text as a comparison baseline. We leveraged the efficiency of single-elimination tournaments, which always require a linear number of matches relative to the number of participants. For brevity, we will refer to single-elimination tournaments simply as tournaments throughout the paper. Varco Arena conducts a tournament for each prompt across all participating LLMs and then computes Elo ratings (Elo and Sloan 1978) based on the results of all matches occurred. By performing tournament matches for multiple LLMs on each prompt, we obtain relative win rates between all possible model pairs within a controlled number of matches. These win rates are then interpreted as Elo ratings, allowing us to create a comprehensive ranking of participating LLMs based on their relative performance.
We validate the effectiveness of our approach, Varco Arena, against the current practice of LLM benchmarks, anchored comparison to reference, through two experiments: one is unbiased judge simulation and the other is running Varco Arena with LLM-as-a-judge on Arena-Hard-Auto benchmark dataset. These experiments confirm that our reference-free benchmarking approach outperforms current practices in terms of rank reliability, given the same number of comparisons. Notably, the simulation experiment, which emulates outcomes under an unbiased judge and controls for factors such as the number of participant LLMs, test prompts, and judge precision, demonstrates that the empirical superiority of our reference-free approach is not coincidental. Instead, it reveals the effectiveness of direct comparison between LLM pairs compared to the current practice of mediated comparison with references for ranking LLMs.
To summarize, our key contributions are as follows:
  1. 1.
    We introduce Varco Arena, a reference-free benchmarking method for LLMs that employs direct comparison of outputs through a single-elimination tournament structure across test prompts.
  2. 2.
    We demonstrate, through conceptual validation and empirical results, that Varco Arena delivers more reliable rankings than current practices of comparing LLM outputs to reference texts, while requiring fewer comparisons.
2 Preliminaries: Quantifying Generation Ability
Quantifying an LLM’s generation ability is crucial for ranking models, but it presents several challenges. The outcomes of comparisons between two LLMs are often probabilistic, influenced by various factors such as test prompts and the subjective nature of human preferences. A straightforward approach is to evaluate an LLM’s performance across diverse test prompts, which approximates its capability in real-world scenarios. Two popular measures for LLM generation ability are: win rate against reference text and elo rating.
2.1 Win rate Against the Reference Text
Several benchmarks quantify response generation quality, with AlpacaEval (Li et al. 2023) and Arena-Hard (Li et al. 2024) being representative examples. These benchmarks employ LLMs as judges for automated evaluation of generation capability. Given a prompt and two responses, the judge LLM is tasked with determining the superior response. These benchmarks include predefined sets of test prompts and reference outputs that serve as quality standards. The evaluation process involves comparing the LLM’s output against the reference to assess response quality. The LLM’s win rate across test prompts is used as a measure of its generation proficiency.
2.2 Elo Rating
The Elo rating system, introduced by Elo and Sloan, has since become a popular method for quantifying performance levels in competitive sports and, more recently, for evaluating the generative abilities of LLMs. The primary purpose of the Elo rating system is to represent a participant’s skill level as a single scalar value, enabling the prediction of relative win rates between participants who have never directly competed. Assessing the superiority of one LLM over another in terms of generative capability shares several similarities with determining winners in competitive sports. Varying test prompts may yield different results, similar to how weather or other factors can influence the outcome of a sports match. While a superior team or LLM is not guaranteed to outperform its inferior counterpart in every instance, it tends to succeed more frequently. With Elo ratings, one can expect the relative chance of winning between a1 pair of players. The expected win rate between a pair of players $i$ and $j$ is computed as follows:
[…]
Computing accurate Elo ratings requires a sufficient number of matches to estimate relative win rates between pairs of participants. The resulting matrix of relative win rates is then used to estimate Elo ratings through logistic regression, similar to multivariate logistic regression with a sigmoid function. In Equation 1, $P(i>j)$ represents the expected win rate of participant $i$ over $j$, $R_{i}$ is the Elo rating of participant $i$, and $\Delta_{\text{ij}}$ is the Elo rating difference between participants $j$ and $i$ ($R_{j}-R_{i}$).
Chatbot Arena presents a leaderboard of LLMs, with Elo ratings computed from matches evaluated by a large user base. Users prompt a pair of LLMs and submit their judgments of which one responded better. Although Chatbot Arena relies on manual evaluation of matches, it offers an intuitive method for comparing LLMs using Elo ratings.
3 Varco Arena
Varco Arena represents our novel approach to reference-free benchmarking of LLMs. Rather than comparing LLM responses to references, we propose iterated tournaments over benchmark dataset to build leaderboards. This addresses concerns about effectiveness and scalability of using direct comparison over traditional benchmarking practice of relying on references. Section 3.1 and Algorithm 1 detail how Varco Arena runs iterated tournaments as we suggest. We then explain our rationale for using direct comparison of LLM responses over bridged comparing to references (Section 3.2), and argue why iterated tournaments can produce effective LLM leaderboards (Section 3.3).
Algorithm 1 Running Iterated Tournaments over Benchmark for multiple LLMs
1:Test prompts $X$, LLMs $M$, outputs $O_{ij}$ = $m_{j}(x_{i})$
2:Ranked LLMs with Elo ratings
3:function Match($m_{1},m_{2},x$)
4:     if $\text{IsBetter}(O_{x,m_{1}},O_{x,m_{2}})$ then
5:         return $m_{1}$
6:     else
7:         return $m_{2}$
8:     end $\triangleright$ omitted byes for conciseness
9:end function
10:function SingleElim(players, $x$, results)
11:     if ||||players|=2|=2| = 2 then
12:         $m_{\text{win}}$leftarrow Match(players[0], players[1], $x$)
13:         results.append($m_{\text{win}}$)
14:         return $m_{\text{win}}$
15:     end if
[…]
17:     l_bracket $\leftarrow$ SingleElim(players[:mid], $x$, results)
18:     r_bracket $\leftarrow$ SingleElim(players[mid:], $x$, results)
19:     return SingleElim(l_bracket + r_bracket, $x$, results)
20:end function
21:function Tournament2Leaderboard($X,M$)
22:     results $\leftarrow$ []
23:     for $x_{i}\in X$ do
24:         SingleElim(Shuffled($M$), $x_{i}$, results)
25:     end for
26:     return ComputeElo(results)
27:end function
3.1 Iterating Tournaments over Benchmark Dataset to Elo ratings
Figure 1 and Algorithm 1 illustrate the methodology of Varco Arena for benchmarking Large Language Models (LLMs) using a tournament approach. In this context, $|X|$ denotes the cardinality of the prompt set in the benchmark dataset (i.e. number of prompts in the dataset). As described running Varco Arena invokes tournament of participant LLMs for each prompt in benchmark dataset, which we will refer to as iterated tournaments.
The application of tournament structures to LLM benchmarking presents both opportunities and challenges. While a single tournament effectively identifies a champion, it does not provide the granularity required for a comprehensive ranking system; it only declares a champion, leaving the precise ordering of other participants ambiguous. To retain the efficient nature of tournaments but to accomplish a fine-grained ranking of LLMs, what modifications should be applied to? To mitigate the aforementioned ranking resolution issue, we propose conducting independent tournaments for each benchmark prompt, with randomized initial match-ups. This approach, which we term iterated tournaments, offers several advantages:
  1. 1.
    It facilitates the resolution of ties among non-champion participants from previous tournaments.
  2. 2.
    It reduces the impact of unfavorable match-ups in any single tournament.
  3. 3.
    While not providing exhaustive pairwise comparisons, the accumulated matches between model pairs enable more precise estimation of relative win rates and Elo ratings, culminating in a well-aligned overall ranking.
In terms of total number of matches, our iterated tournaments approach requires $|X|*(n_{\text{models}}-1)$ which is always $|X|$ numbers of comparisons saved compared to anchored comparison to references. A favorable trait of conducting tournaments for each prompt is that every participating LLM engages in at least one competition per prompt. The number of matches an LLM faces per test prompt varies from 1 to $\lceil\text{log}_{2}(n_{\text{models}})\rceil$, based on its performance. Allocating more matches for better LLMs will help precise separation between higher rated models.
We further elaborate and argue why aggregating multiple tournaments should reach a good sorting of participant LLMs in Section 3.3. Additionally, we provide prospect match numbers LLMs would confront within, offering a comprehensive view of the method’s efficiency and effectiveness.
3.2 Comparing to Reference is not Always Helpful
Here, we inspect into failure modes of relying on references for ranking LLMs, which further elaborates the previously mentioned concerns on current practices for LLM benchmarking. Consider a highly accurate judge capable of definitively comparing two outputs of different quality without errors or ties. In this scenario, all matches involving participant LLMs contribute to their relative win rates and, ultimately, to their final Elo ratings (Section 3.3 and Equation 2). Consequently, all matches involving the LLM of interest contribute to Elo rating and ranking of the LLM.
When evaluating LLM performance relative to reference texts, even with the same precise judge, the comparison doesn’t always provide a reliable measure of relative win rates. Suppose we have only two participants of interest $\text{M}_{1}$ and $\text{M}_{2}$ and an instruction $X_{i}$ to prompt these LLMs. There might be two corresponding responses $\text{M}_{1}(X_{i})$, $\text{M}_{2}(X_{i})$ and also a reference text $Y_{i}$. Equation 2 illustrates four possible scenarios. In two of these cases, the match outcome is useful for comparing two LLMs, specifically when the reference text loses to one model and wins against the other, working as a useful anchor for comparison. In the other two cases, comparing to $Y_{i}$ fails to differentiate between $\text{M}_{1}(X_{i})$ and $\text{M}_{2}(X_{i})$.
$\begin{array}[]{l}\text{M}_{1}(X_{i})\\ ~{}~{}~{}~{}~{}\text{vs.}~{}~{}~{}~{}~{}\rightarrow\\ \text{M}_{2}(X_{i})\end{array}\begin{cases}\text{M}_{1}(X_{i})>Y_{i}>\text{M}_{2}(X_{i})&\\ ~{}~{}~{}~{}~{}\text{or}&\text{useful}\\ \text{M}_{1}(X_{i})<Y_{i}<\text{M}_{2}(X_{i})&\\ &\\ \text{M}_{1}(X_{i}),~{}\text{M}_{2}(X_{i})>Y_{i}&\\ ~{}~{}~{}~{}~{}\text{or}&\text{useless}\\ \text{M}_{1}(X_{i}),~{}\text{M}_{2}(X_{i})<Y_{i}&\\ \end{cases}$ (2)
This analysis reveals a potential weakness in benchmarks that rely on anchored comparisons to reference texts. If the reference texts are not carefully calibrated in terms of difficulty, the ranking of LLMs could be determined by only a small portion of the test prompts. Specifically, if reference texts are consistently superior or inferior to most model outputs, a large number of comparisons become uninformative. This issue is mitigated when directly comparing model outputs without a reference text, suggesting that direct comparisons may provide a more robust method for ranking LLMs.
3.3 Iterated Tournaments as a Rough Approximation of Merge Sort
In this section, we argue that iterated tournaments will reach to a good sorting of participants LLMs comparing the similarities in merge sort algorithm and iterated tournament regarding the presumptions of Elo model for LLM capability.
Our approach is predicated on a key assumption: although the prompts in the benchmark dataset are diverse, they can be considered approximately equivalent in their capacity to differentiate LLM performance. This assumption of ‘indistinguishable prompts’ aligns with the Elo model of LLM capability, which posits that prompt-specific variations should not significantly affect the expected outcome likelihood of a single match. [^1] Given the simplicity of the Elo model’s assumptions regarding LLM capability, we posit that ranking LLMs with Elo ratings for a mixture of disparate capabilities may be inappropriate, as a higher Elo rate does not guarantee uniform mastery across multiple capabilities.
Building on this assumption, we propose that conducting tournaments across the benchmark prompts can lead to an accurate ranking of LLMs. We suggest that single-elimination tournaments function analogously to merge sort, with a crucial distinction: they perform only the initial comparison at each step of a merge operation and discard the losing side (i.e. only the winner proceeds to the next rounds). This makes a single tournament perform $n-1$ comparisons in total. This number of comparisons is less than $\mathcal{O}(n\text{log}_{2}n)$ which most of the efficient sorting algorithms requires. However, regarding the assumption of ‘indistinguishable prompt’ we mentioned above, number of tournament matches from the other dataset prompt would help filling the lack of comparion matches. By iterating over the benchmark dataset with randomized initial match-ups, these tournament matches can eventually approximate the complete set of comparisons required for merge sort, given a sufficient number of benchmark prompts ($|X|$).
Our approach provides $\mathcal{O}(|X|$/$\text{log}_{2}n_{\text{model}})$ times more comparisons [^2] Where $\mathcal{O}(|X|$/$\text{log}_{2}n_{\text{model}})=\frac{|X|(n_{\text{model}}-1)}{\mathcal{O}(n_{\text{model}}\text{log}_{2}n_{\text{model}})}$ representing the ratio between matches in iterated tournaments and matches needed for sorting LLMs with an ideal judge. than the minimum required for merge sorting $n_{\text{model}}$ LLMs. In practice, number of matches to figure which model is better from a pair is much larger since the judge model will not be ideal and also performance of LLMs will be stochastic regarding the prompt it was asked to answer. Nevertheless, the large number of prompts ($|X|$), typically ranging from hundreds to thousands, lead us to the worth of empirical validation of the idea, that iterated tournaments will accurately sort LLMs with confidence. This foundational argument underpins our unbiased judge simulation experiment to test the potential of tournament approach (Varco Arena).
4 Experiments
We run two experiments with different settings for comparing tournament and anchored comparison approaches. The first experiment (Section 4.2) simulates an unbiased judge to control various factors that affect LLM benchmarking. This simulation tests our foundational propositions for Varco Arena design (mentioned in Section 3.2 and 3.3) under a more controlled, simplified environment. It eliminates noisy factors such as potential biases of LLM judges (Park et al. 2024).
In subsequent empirical experiments (Section 4.3), we use gpt-4o[-mini] as a judge, to validate the effectiveness of our tournament approach against current LLM benchmarking practices. By presenting both simulation and empirical results, we aim to demonstrate the effectiveness of our tournament approach. We first describe the common experimental settings before getting into specific details in the following subsections.
4.1 Chatbot Arena Leaderboard Ratings as Ground-Truth LLM Rankings
We compare the results of each benchmarking approach against the rankings from the Chatbot Arena leaderboard. Chatbot Arena is currently considered one of the most reliable leaderboards due to its large number of human preference annotations. In this platform, users are presented with responses from a pair of LLMs to the same query and asked to choose which response they prefer. Given the substantial number of votes and the wide range of submitted prompts used for model comparison, the resulting ranks from Chatbot Arena are likely accurate enough to serve as ground-truth rankings.
4.2 Experiment 1: Simulating an Unbiased Judge
We designed a simple stochastic modeling experiment to simulate LLM matches following the Elo preference model. Our model assumes that the primary factor governing the outcome of each match is the Elo rating gap ($\Delta_{\text{ij}}$), as shown in Equation 3. We also introduced a secondary factor, judge precision ($P_{\text{judge}}$), which represents the likelihood that a judge correctly declares the higher-rated LLM as the winner. The observed outcome of a single match is sampled according to the likelihood of model $i$ winning against model $j$ ($P_{\text{predict}}(i>j)$). This probability is the product of the judge precision and the estimated relative win rate expected by ground-truth Elo ratings ($P_{\text{gt}}(i>j)$). This simplified simulation model aligns with the fundamental principle of the Elo system: the outcome of a match between a pair of models is primarily determined by their skill gap ($\Delta_{\text{ij}}$). This gap is derived from past wins and losses, effectively condensing all sophisticated factors into a single mean value.
$P_{\text{predict}}(i>j)=P_{\text{judge}}*P_{\text{gt}}(i>j)=P_{\text{judge}}*\frac{1}{1+10^{\Delta_{\text{ij}}/400}}$ (3)
The details of our simulation settings are as follows:
Ground-truth Elo ratings: We extracted Elo ratings from the English category of Chatbot Arena as of June 23, which includes one million user-submitted prompts and judgments (approximately 60% of the total submissions to the platform). While any set of Elo ratings could be valid for this simulation, we chose to use real values computed from human preferences.
Judge Precision: In real-world, judge precision would be an adaptive value that varies based on the prompt-response pair being judged or the method of prompting the LLM-as-a-judge. In our experiments, we controlled judge precision, ranging from 0.6 to 0.9.
Number of participant LLMs ($n_{\text{model}}$) and benchmark dataset size ($|X|$): We varied these factors to assess the robustness of tournament and anchored comparison approaches under diverse conditions. A larger number of participants may require more matches to achieve precise results. Similarly, the benchmarking environment may be data-poor or data-rich. We emulated the varying dataset size by sampling different numbers of matches from $P_{\text{predict}}$.
The simulation process follows these steps:
  1. 1.
    Select participants from the top of the Chatbot Arena leaderboard, avoiding ties (e.g., selecting only one model from a tied rank).
  2. 2.
    Compute expected relative win rates ($P_{\text{gt}}$, Equation 3) from the participants’ Elo ratings.
  3. 3.
    Sample match outcomes across LLM pairs for tournaments or anchored comparisons. An unbiased judge with precision $P_{\text{judge}}$ (Equation 3) determines the winner according to Equation 3, considering only the Elo gap $\Delta_{\text{ij}}$.
  4. 4.
    Repeat step 3 for the designated benchmark dataset size $|X|$.
  5. 5.
    Calculate scores for ranking:
    1. (a)
      For anchored comparison, use the win rate against the reference LLM (we choose gpt-4-1106-preview which records 1233 in Elo rate).
    2. (b)
      For tournaments, compute Elo ratings from the recorded match outcomes from every tournaments performed.
  6. 6.
    Rank models based on these scores.
  7. 7.
    Compute the Spearman correlation between the simulated rankings (6) and the ground-truth rankings (1).
We perform 50 trials per simulation configuration to mitigate randomness from tournament brackets and sampling.
4.3 Experiment 2: Running Varco Arena using gpt-4o[-mini] as Judge
We conducted an empricial experiemnt of Varco Arena to validate our proposed method. We conducted tests for both tournament approach and anchored comparison approach using 20 top-ranked models from the Chatbot Arena leaderboard. This experiment differs from our modeling simulation in several key aspects. We elaborate on the details of the experiments in order of how we prompted the LLM judges and our practice to managing API costs while conducting a large number of experiments.
Dataset for Testing Benchmarking Approaches
In order to test the benchmarking approaches, tournament and anchored comparison, we need two sorts of data. One is a prompt to be asked for LLMs and the other is a reponse by each LLM given the prompt. For the benchmark dataset, we choose Arena-Hard-Auto (Li et al. 2024). Arena-Hard-Auto benchmark prompts are carefully assorted from Chatbot Arena which are user-typed queries to LLMs. The total number of prompts in Arena-Hard-Auto is 500, each 2 instances for 250 subtopics. Choosing Arena-Hard-Auto as our benchmark dataset is natural since it minimizes the disparity between our benchmarking environment and the ground truth environment where human annotators actually judged LLMs. As suggested in Arena-Hard-Auto, we use the responses of gpt-4-0314 as reference outputs. For responses of the LLMs to be ranked, we use the reserved outputs of the top 21 models from Huggingface space of the Arena-Hard-Auto Browser. [^3] Extracted from Jul 6 commit (fd42026)
Participant LLMs
As a list of models to be ranked, we choose 20 LLMs from the top of ChatBot Arena leaderboard in Hard prompts category which is considered to synchronize the most to Arena-Hard-Auto. To avoid egocentric bias, we exclude judge models from the participant LLMs. We use gpt-4o-2024-05-13 and *4o-mini as our judge, gpt-4-0314 as a reference (anchor) model to perform anchored comparisons on. You can see the evaluation prompt on Listing 1 at Appendix.
LLM Judges
We deploy gpt-4o and gpt-4o-mini as judges. To prompt the pairwise comparison of the responses from the participant LLMs, we refer to (Zeng et al. 2024), equally for both tournament and anchored comparison approaches. To avoid position bias (Wu and Aji 2023), we alternate the position of the model responses for prompting evaluation. Further details on the LLM-as-a-judge configuration are provided in Appendix A.5. While using gpt-4o-2024-05-13 as a judge provides a realistic setting for LLM benchmarking, we also explore gpt-4o-mini-2024-07-18 to accommodate users with budget constraints.
Assuring Statistical Significance of the Results within Budget
To ensure a statistically significant number of trials for each experiment while staying within budget, we utilize OpenAI’s Batch API to prepare full-grid match outcomes (i.e., all-play-all matches for every prompt) in a cache file, allowing us to reuse these outcomes. Each empirical experiment consists of 500 trials per setting, with results represented using whisker plots or summary statistics such as median values. When experimenting with a subset of the Arena-Hard-Auto benchmark ($|X|<500$), we sample a stratified subset of the benchmark dataset for each new trial.
To summarize the differences in settings: Experiment 1 (simulating unbiased judge) uses ground truth Elo ratings from Chatbot Arena (English category) and control parameters of our choices (e.g., $P_{\text{judge}}$, $|X|$, and $n_{\text{model}}$) to predict LLM rankings, which are then compared to the ground-truth rankings. Experiment 2 (running Varco Arena with LLM judges) benchmarks the top 20 models from Chatbot Arena (Hard prompts category) using the Arena-Hard-Auto benchmark dataset. For both tournament and anchored comparison approaches, we employ Spearman correlation as a correlation measurement to see how well the benchmarking approaches aligns to the ground-truth leaderboard.
5 Results and Discussion
We report on the reliability of Varco Arena for simultaneously ranking multiple LLMs. Both simulated and empirical results demonstrate that the tournament approach of Varco Arena produces ranks more closely aligned with ground-truth Elo ranks.
5.1 Experiment 1: Unbiased Judge Simulation
Figure 2: Simulated results comparing tournament and anchored comparison approaches with an unbiased judge. The tournament approach consistently demonstrates superiority in terms of Spearman correlation across various control variables: number of participant LLMs ($n_{\text{models}}$), number of prompts in the benchmark set ($|X|$), and judge precision ($P_{\text{judge}}$).
Figure 2 reveals distinct gaps in Spearman correlation, indicating that the tournament approach demonstrates greater reliability compared to the anchored comparison approach. These significant gaps persist across various controlled parameters, including the number of participating LLMs ($n_{\text{models}}$), prompts in the benchmark set ($|X|$), and judge precision ($P_{\text{judge}}$). While the simulation necessarily abstracts real-world complexity, the consistent performance patterns observed under systematically varied conditions align with the empirical findings from Experiment 2, suggesting potential robustness of the tournament approach beyond the specific conditions tested empirically.
5.2 Experiment 2: Running Varco Arena with gpt-4o[-mini] Judge
Figure 3: Tournament and anchored comparison approach run with gpt-4o (left) and gpt-4o-mini (right) judge. Tournament approach constantly records higher Spearman correlation across the varying number of benchmark prompts used. Results summary is on Table 1.
Our empirical experiments corroborate the trends observed in the simulation. As illustrated in Figure 3 and Table 1, while the gaps are less pronounced compared to the simulation results (Figure 2), the Varco Arena results obtained through the tournament approach consistently demonstrate higher correlation across varying benchmark sizes ($|X|$, denoted as $n_{\text{prompts}}$ in Figure 3). This empirical evidence further strengthens our argument for the effectiveness of the tournament approach.
Spearman corr. ($\uparrow$) $|X|$ = 50 100 250 475 500
anchored (4o) 0.895 0.935 0.963 0.966 0.964
tournament (4o) 0.905 0.940 0.960 0.970 0.970
anchored (4o-mini) 0.895 0.908 0.917 0.916 0.912
tournament (4o-mini) 0.901 0.919 0.931 0.933 0.933
Table 1: Spearman correlation ($\uparrow$) varying over size of the benchmark set ($|X|$) for each benchmarking approach. Values shown are median from 500 trials. See Figure 3 for detailed whisker plots.
Varco Arena (labeled as tournament) consistently demonstrates higher correlation with the ground-truth leaderboard, regardless of deploying gpt-4o or gpt-4o-mini as the judge. Using the less expensive gpt-4o-mini as the judge did not compromise much in terms of correlation compared to its costlier counterpart. Notably, this tournament approach achieves comparable or superior correlation without the need for reference texts, while also reducing the number of required comparisons for benchmarking by $|X|$, further saving costs. The advantage of benchmarking by tournaments becomes more pronounced as the benchmark dataset size, $|X|$, increases.
5.3 Incorporating a New LLM into an Existing Leaderboard
While our main focus has been on ranking multiple LLMs simultaneously, it’s worth considering how our approach might apply to the common scenario of adding a single new model to an existing leaderboard. We explored a binary search-like placement method for this purpose (detailed in Appendix 2), but found it yielded less stable results than the reference output approach (Table 2). We attribute this instability to two factors: (1) The reduced number of comparisons (only $|X|$ comparisons per model, the minimum for a tournament run). (2) Incomplete coverage of the benchmark dataset’s prompts.
However, the concept of a reference model as an anchor could still be valuable in this context. One could estimate the win rate against a anchor of one’s own choice for all models in the leaderboard, using this as a ranking score. For leaderboards built with Varco Arena, any missing matches could be imputed using expected relative win rates derived from Elo ratings or directly from existing match records. This approach offers a potential compromise between the stability of reference-based methods and the comparative advantages of our tournament approach.
$\bar{|\Delta_{\text{rank}}|}$ ($\downarrow$) gt=1-6 7-13 14-19 (20) total avg.
binary search (4o) 0.92 1.84 2.13 1.72
anchored (4o) 1.98 1.55 1.57 1.39
binary search (4o-mini) 1.27 1.82 1.21 1.5
anchored (4o-mini) 1.00 1.43 1.43 1.37
Table 2: Binary search vs. Anchored comparison for the insertion of a new LLM: Mean rank deviation ($\bar{|\Delta_{\text{rank}}|}$) from ground-truth leaderboard. Result of binary search placement and anchored comparison insert by gpt-4o[-mini] judge are provided with bootstrapped 95% confidence interval (500 trials, 1000 samples, $|X|$=500, Arena-Hard-Auto).
6 Related Works
6.1 Elo Modeling for LLM Benchmarking
Utilizing Elo modeling of LLM capability based on human preference is a recently recognized means for benchmarking (Boubdir et al. 2023). Chatbot Arena (Chiang et al. 2024) is introduced as a benchmark platform that targets live benchmarking of LLMs toward open-ended generation tasks with the proper use of Elo ratings. Except that the platform requires massive human judge annotation, Chatbot Arena is considered a successful example of dynamic benchmarking for open-ended tasks as suggested in (Kiela et al. 2021). RAGElo (Rackauckas, Câmara, and Zavrel 2024) deploy Elo model with LLM as a judge (Zheng et al. 2023) to improve retrieval augmented generation (RAG) pipeline which indicates that modeling LLM capability with Elo rating is practically effective.
6.2 Reference-free Evaluation
Alleviating the need for reference text has long been considered an attractive goal in Natural Language Generation (NLG) evaluation. Recent advances in LLMs have shown promising results in their ability to assess open-ended responses (Jauhiainen and Guerra 2024). Also, findings from the WMT23 metrics task in neural machine translation (Freitag et al. 2023) have demonstrated that reference-free quality estimation metrics outperform reference-based metrics in terms of correlation to human evaluation. Specifically, metrics such as XComet (Guerreiro et al. 2023) and MetricX (Juraska et al. 2023) have shown superior performance, particularly when dealing with scenarios involving poor-quality references.
7 Conclusion
We proposed Varco Arena, a reference-free LLM benchmarking approach for measuring generative capabilities of LLMs. Varco Arena deploys a tournament-style approach with direct pairwise comparisons, eliminating the need for reference outputs. Our simulation experiments demonstrate that this tournament approach achieves aligned LLM rankings better than current practice of reference use, under unbiased judging conditions. We also report the empirical effectiveness of Varco Arena using gpt-4o[-mini] as a judge, which reaffirms the efficacy of our approach in practice by showing improved correlation with Chatbot Arena leaderboard rankings. Through extensive experimentation and rigorous statistical analysis, we strongly suggest that our reference-free tournament approach, Varco Arena, provides a superior solution for LLM benchmarking, requiring fewer computational resources.
Appendix A Appendix
A.1 Machine Requirements for Experiments
All the experiments we performed are conducted by querying API and post-processing those, which does not require GPU machine locally. Experiments could be run on personal desktops. The lowest specification of the machine we deployed had i5-8400 CPU, 16 GiB RAM.
Figure 4: Elo ratings of the model with gpt-4o judge on the full set of Arena-Hard-Auto (Li et al. 2024) prompts. Varco Arena result (bootstrapped median over 1000 samples of 500 trials) is in blue, plotted alongside the ratings from the ground truth leaderboard in red (Chatbot Arena, Hard prompts category). Error bars are 95% confidence intervals.
A.2 Elo ratings from Varco Arena compared to Human Annotations
Figure 4 shows the Elo ratings computed out of Varco Arena. For judge, we used gpt-4o. As mentioned in the caption, the Elo ratings are bootstrapped median value from 500 trials. 95% confidence intervals also plotted as an error bar, which look negligible in scale compared to observed values. Matches are performed over Arena-Hard-Auto benchmark dataset (500 prompts).
A.3 Binary search vs. Win rate over reference
Algorithm 2 Binary Search for Enlisting new LLM to a leaderboard
1:Leaderboard $L$, new model $m_{\text{new}}$, test prompts $X$, outputs $O_{ij}$, assumes $|X|>|L|>n_{\text{comparisons}}$
2:Updated leaderboard $L^{\prime}$ with $m_{\text{new}}$ placed
3:$n_{\text{comparisons}}\leftarrow\lfloor\log_{2}(|L|)\rfloor$
4:$n_{\text{matches}}\leftarrow\lfloor|X|/n_{\text{comparisons}}\rfloor$
5:function BinarySearchPlacement($L,m_{\text{new}}$)
6:     X $\leftarrow$ Shuffle(X)
7:     X $\leftarrow$ concat(X;X)
8:     low $\leftarrow 0$
9:     high $\leftarrow|L|-1$
10:     while low $\leq$ high do
11:         mid $\leftarrow\lfloor(\text{low}+\text{high})/2\rfloor$
12:         wins $\leftarrow 0$
13:         for $i\leftarrow 1$ to $n_{\text{matches}}$ do
14:              $x\leftarrow$ $X$.pop()
[…]
16:                  wins $\leftarrow$ wins $+1$
17:              end if
18:         end for
19:         if wins $>n_{\text{matches}}/2$ then
20:              high $\leftarrow$ mid $-1$
21:         else if wins $<n_{\text{matches}}/2$ then
22:              low $\leftarrow$ mid $+1$
23:         else if $|X|>$then
24:              continue $\triangleright$ Ensure tie
25:         else
26:              return mid, tie $\triangleright$ Tie
27:         end if
28:     end while
29:     return low, non-tie $\triangleright$ Position found
30:end function
31:function UpdateLeaderboard($L,m_{\text{new}}$)
32:     position, istie $\leftarrow$ BinarySearchPlacement($L,m_{\text{new}}$)
33:     $L^{\prime}\leftarrow L$.insert(position, $m_{\text{new}}$, istie)
34:     return $L^{\prime}$
35:end function
Binary Search
We tried binary search placement of a newly added LLM to the leaderboard without reference text in Table 3. Details of how we implemented binary search are attached in Appendix 2. It turns out that binary search based on leaderboard ranks is not as reliable as the current approach of scoring the newcomer to the reference outputs. The number of judge operations performed is equivalent to the matches allocated to the least-performant model in a tournament, which is $|X|$ (i.e. maximum possible matches that an LLM could have is $|X|*\text{log}_{2}n_{\text{model}}$). Within the size of the benchmark prompts ($|X|$), binary search is incompatible with the current approach of using reference instead.
Converting Elo Table back to Win Rate
Assuming we preserved a set of match results and model outputs from the last benchmarking, we could benefit from those to perform insertion. One could pick an appropriate anchor LLM as a reference in a leaderboard to estimate the skill of a newcomer. Using previous matches from the tournaments that built the leaderboard could be used for estimating win rates over the reference. This is the same as converting the Elo table into a win rate leaderboard. Since the leaderboard is not built with full-grid matches but with tournaments, there would be some missing matches against the reference regardless we have picked. There are two ways to estimate the win rate over the reference model. We could just count the matches given are enough in amount, or we could also convert Elo ratings back to $P(i>a)$ to use it directly for scoring for the model ranks in the leaderboard. Reminding that Elo rating is purposed for expecting a likely outcome of the match, this should work. After this win rate of the newcomer model $P^{*}(n>a)=\frac{\text{count(n wins)}}{|X|}$ could be directly compared for enlisting.
$|\Delta_{\text{rank}}|$ ($\downarrow$) gt=1 2 3 4 5 6 avg.
binary search 0.09 1.24 1.75 1.55 1.26 1.10 0.92
(4o) (.04/-.03) (.14/-.14) (.09/-.09) (.07/-.06) (.08/-.08) (.10/-.09)
anchored 0.00 1.01 1.95 2.00 0.96 0.30 1.98
(4o) (0.00/0.00) (0.01/-0.01) (0.02/-0.02) (0.00/0.00) (0.02/-0.02) (0.04/-0.04)
binary search 0.52 0.85 0.59 2.03 1.20 2.45 1.27
(4o-mini) (.09/-.07) (.12/-.11) (.10/-.09) (.02/-.02) (.05/-.05) (.07/-.06)
anchored 0.00 0.00 1.00 2.00 2.00 1.00 1.00
(4o-mini) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00)
7 8 9 10 11 12 13 avg.
1.31 1.27 2.22 1.74 2.27 2.23 1.86 1.84
(.10/-.10) (.11/-.11) (.14/-.12) (.09/-.09) (.12/-.11) (.12/-.12) (.07/-.07)
0.30 3.68 1.09 1.03 2.97 0.78 1.00 1.55
(0.04/-0.04) (0.04/-0.04) (0.03/-0.03) (0.02/-0.01) (0.02/-0.02) (0.05/-0.05) (0.00/0.00)
0.69 0.85 3.89 1.95 2.10 2.37 0.88 1.82
(.07/-.06) (.09/-.09) (.12/-.11) (.06/-.05) (.03/-.03) (.10/-.11) (.12/-.11)
0.51 0.52 3.50 1.00 1.00 3.00 0.50 1.43
(0.49/-0.51) (0.48/-0.52) (0.49/-0.51) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.50/-0.50)
14 15 16 17 18 19 20 avg.
1.40 3.07 0.80 1.47 5.00 0.96 - 2.13
(.04/-.05) (.11/-.11) (.08/-.09) (.05/-.04) (.11/-.11) (.08/-.09)
2.00 2.00 1.00 1.21 3.00 0.21 - 1.57
(0.00/0.00) (0.00/0.00) (0.00/0.00) (0.03/-0.04) (0.00/0.00) (0.04/-0.03)
1.45 4.20 0.19 0.08 1.09 1.08 0.40 1.21
(.07/-.08) (.17/-.17) (.07/-.06) (.03/-.02) (.05/-.05) (.05/-.05) (.07/-.07)
1.00 2.00 2.00 1.00 1.00 3.00 0.00 1.43
(0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00) (0.00/0.00)
Table 3: Binary search vs. Anchored comparison: Mean rank deviation ($|\Delta_{\text{rank}}|$) from ground-truth leaderboard. Result of binary search placement and anchored comparison insert by gpt-4o[-mini] judge are provided with bootstrapped 95% confidence interval (500 trials, 1000 samples, $|X|$=500, Arena-Hard-Auto (Li et al. 2024)).
A.4 Separability In terms of Confidence Interval
To see how well the two benchmarking approach (anchored comparison and tournament approach) separates LLMs in adjacent ranks, we provide scatter plot of Elo rating and win rate paired with error bar (95% confidence interval). We present the both results of using gpt-4o (Figure 5) and gpt-4o-mini (Figure 5) as a judge. Inside the each plot, inseparables indicates the cases where any pair of datapoint co-includes each other within their range of error bars, and overlap means a certain datapoint is within some other’s range of error, when it is one-sided.
Figure 5: gpt-4o result of anchored comparison and tournament approach. 1000 bootstrapped median from 500 observations used for confidence interval estimation.
Figure 6: gpt-4o result of anchored comparison and tournament approach. 1000 bootstrapped median from 500 observations used for confidence interval estimation.
A.5 Judge configuration
Evaluation Prompt
We use the prompt from LLMBar. The prompt depicted in Figure 1. We added 4 questions for criteria of our own to Metrics.txt prompt of (Zeng et al. 2024). You can refer to the original prompt in LLMBar github.
Decoding Parameters
We did not configure decoding parameters of judge LLMs (gpt-4o[-mini]), which its temperature defaults to 1. The only parameter we have adjusted is maximum number of tokens to be generated, which for our prompt is less than 6 (i.e. The output of our prompt is (a) or (b)). To avoid position bias, we alternated the position of the responses from a certain model across the benchmark prompt.
Listing 1: LLMBar prompt (Metrics.txt)
1PROMPTS = [
2 {
3 role”: system”,
4 content”: ”””You are a helpful assistant in evaluating the quality of the outputs for a given instruction. Your goal is to select the best output for the given instruction.”””,
5 },
6 {
7 role”: user”,
8 content”: ”””Select the Output (a) or Output (b) that is better for the given instruction. The two outputs are generated by two different AI chatbots respectively.
9
10Here are some rules of the evaluation:
11(1) You should prioritize evaluating whether the output honestly/precisely/closely executes the instruction, then consider its helpfulness, accuracy, level of detail, harmlessness, etc.
12(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
13(3) You should avoid any potential bias and your judgment should be as objective as possible. For example, the order in which the outputs were presented should NOT affect your judgment, as Output (a) and Output (b) are **equally likely** to be the better.
14
15Do NOT provide any explanation for your choice.
16Do NOT say both / neither are good.
17You should answer using ONLY Output (a)” or Output (b)”. Do NOT output any other words.
18
19# Instruction:
20{instruction}
21
22# Output (a):
23{response_a}
24
25# Output (b):
26{response_b}
27
28# Questions about Outputs:
29Here are at most three questions about the outputs, which are presented from most important to least important. You can do the evaluation based on thinking about all the questions.
30* Does the output well satisfy the intent of the user request?
31* If applicable, is the output well-grounded in the given context information?
32* Does the output itself satisfy the requirements of good writing in terms of:
33 1) Coherence
34 2) Logicality
35 3) Plausibility
36 4) Interestingness
37
38
39# Which is better, Output (a) or Output (b)? Your response should be either Output (a)” or Output (b)”:”””,
40 },
41]
Appendix B Reproducibility Checklist
  • Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes)
  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes)
  • Provides well marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper (yes)
  • Does this paper make theoretical contributions? (yes)
  • All assumptions and restrictions are stated clearly and formally. (yes)
  • All novel claims are stated formally (e.g., in theorem statements). (partial)
  • Proofs of all novel claims are included. (partial)
  • Proof sketches or intuitions are given for complex and/or novel results. (yes)
  • Appropriate citations to theoretical tools used are given. (yes)
  • All theoretical claims are demonstrated empirically to hold. (yes)
  • All experimental code used to eliminate or disprove claims is included. (no)
  • Does this paper rely on one or more datasets? (yes)
  • A motivation is given for why the experiments are conducted on the selected datasets (yes)
  • All novel datasets introduced in this paper are included in a data appendix. (NA)
  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes)
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. (yes)
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. (yes)
  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. (NA)
  • Does this paper include computational experiments? (yes)
  • Any code required for pre-processing data is included in the appendix. (no)
  • All source code required for conducting and analyzing the experiments is included in a code appendix. (no)
  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes)
  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (partial)
  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. (yes)
  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. (partial)
  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. (yes)
  • This paper states the number of algorithm runs used to compute each reported result. (yes)
  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. (yes)
  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). (yes)
  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes)
  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes)

### v2

@@
 Abstract
-The rapid advancement of Large Language Models (LLMs) necessitates robust evaluation methodologies.
-Current benchmarking approaches often rely on comparing model outputs against predefined prompts and reference outputs.
-Relying on predefined reference outputs hinders flexible adaptation of benchmarks to the rapidly evolving capabilities of LLMs.
-This limitation necessitates periodic efforts to prepare new benchmarks.
-To keep pace with rapidly evolving LLM capabilities, we propose a more flexible benchmarking approach.
-Our method, Varco Arena, provides reference-free benchmarking of LLMs in tournament style.
-Varco Arena directly compares LLM outputs across a diverse set of prompts, determining model rankings through a single-elimination tournament structure.
-This direct pairwise comparison offers two key advantages: (1) Direct comparison, unmediated by reference text, more effectively orders competing LLMs, resulting in more reliable rankings, and (2) reference-free approach to benchmarking adds flexibility in updating benchmark prompts by eliminating the need for quality references.
-Our empirical results, supported by simulation experiments, demonstrate that the Varco Arena tournament approach aligns better with the current Elo model for benchmarking LLMs.
-The alignment is measured in terms of Spearman correlation, showing improvement over current practice of benchmarking that use reference outputs as comparison anchors.
+Most existing benchmarking approaches for evaluating the output quality of large language models (LLMs) rely on comparing LLM responses to predefined references.
+Such methods, based on static datasets, quickly become outdated as LLM capabilities and use cases evolve.
+In this work, we introduce VARCO Arena—a novel, cost-effective, and robust benchmarking approach that leverages a single-elimination tournament structure to minimize the number of required comparisons while eliminating the need for static references or costly human annotations.
+We validate our approach through two experiments: (i) a simulation study that examines its robustness under various conditions, and (ii) an empirical evaluation using publicly available benchmark prompts.
+In both experiments, VARCO Arena consistently outperforms current LLM benchmarking practices, achieving stronger correlations with human-established Elo ratings.
+Our results demonstrate that VARCO Arena not only produces reliable LLM rankings but also provides a scalable, adaptable solution for qualitative evaluation across diverse, customized use cases.
+We release our demo and code at [URL placeholder].
+VARCO Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models
+Seonil Son1, Ju-Min Oh1thanks: Current institution: Samsung AI Research, Heegon Jin1thanks: Current institution: Global AI Platform, Cheolhun Jang1, Jeongbeom Jeong1, Kuntae Kim1 1NCSOFT Correspondence: deftson@ncsoft.com
 1 Introduction
 The versatility of Large Language Models (LLMs) stems from their generative capacity to address a wide array of tasks.
-The multi-faceted capability of LLMs enables flexible applications across numerous user scenarios, exemplified by generalist assistants (Ouyang et al.
-2022; Köpf et al.
-2024; Roziere et al.
-2023).
+The multi-faceted capability of LLMs enables flexible applications across numerous user scenarios Ouyang et al.
+(2022); Köpf et al.
+(2024); Roziere et al.
+(2023).
 As versatility emerges as a core attribute of LLMs, the challenge of accurately gauging their skill becomes increasingly significant.
-In response to this challenge, numerous benchmarks evaluating LLM capability have emerged (Hendrycks et al.
-2020; Srivastava et al.
-2023; Zhong et al.
-2024).
+In response to this challenge, numerous benchmarks evaluating LLM capabilities have emerged Hendrycks et al.
+(2020); Srivastava et al.
+(2023); Zhong et al.
+(2024).
 Many LLM benchmarks employ formats amenable to automated scoring.
-Examples include benchmarks for arithmetic problems (Gao et al.
-2022; Patel, Bhattamishra, and Goyal 2021), multiple-choice questions (Lin, Hilton, and Evans 2022), and code execution (Austin et al.
-2021; Chen et al.
-2021).
+Examples include benchmarks for arithmetic problems Gao et al.
+(2022); Patel et al.
+(2021), multiple-choice questions Lin et al.
+(2022), and code execution Austin et al.
+(2021); Chen et al.
+(2021).
 While these benchmarks are valuable, they primarily assess problem-solving abilities.
-The need for LLM benchmark extends beyond this, as the primary value of LLMs lies in their generative capabilities.
-Researchers have primarily relied on pariwise comparison to evaluate and rank the generative capabilities of LLMs.
-A representative example of this approach is Chatbot Arena (Chiang et al.
-2024), which computes Elo ratings based on a massive number of human-annotated pairwise comparisons.
-While Chatbot Arena benefited from the reliability of open-ended prompts from users, acquiring such a large number of annotations for benchmarking LLMs is time-consuming and resource-intensive.
-To address these challenges, several benchmark datasets for generation tasks have been developed.
-These leverage reference responses for ranking LLMs using automated LLM judges (Zheng et al.
-2024) includes AlpacaEval (Li et al.
-2023), Arena-Hard-Auto (Li et al.
-2024), and MTBench (Zheng et al.
-2023).
-Relying on reference outputs offers two key advantages: (1) The number of comparisons required for ranking multiple LLMs reduces linear to the number of LLMs.
-(2) Reference texts establish a quality bar for evaluating responses.
-However, we propose that directly comparing the responses of LLMs, rather than relying on reference outputs, is a more effective way to achieve accurate LLM rankings.
-We suggest that combined with tournament-styled match making, direct comparison of outputs can achieve a better correlation to human preference within the same number of comparisons.
-Additionally, reference-free benchmarking mitigates potential biases in open-ended generation tasks and provides more flexibility compared to static benchmark datasets.
-Figure 1: Varco Arena directly compares LLM output pairs in single-elimination tournament structure rather than comparing LLM outputs to reference text.
-In terms of deciding whether a certain LLM is better or worse compared to the other one, we suggest direct comparison is more intuitive and results in better separability.
-We devised Varco Arena, which directly compares LLM outputs without the need for reference text as a comparison baseline.
-We leveraged the efficiency of single-elimination tournaments, which always require a linear number of matches relative to the number of participants.
-For brevity, we will refer to single-elimination tournaments simply as tournaments throughout the paper.
-Varco Arena conducts a tournament for each prompt across all participating LLMs and then computes Elo ratings (Elo and Sloan 1978) based on the results of all matches occurred.
-By performing tournament matches for multiple LLMs on each prompt, we obtain relative win rates between all possible model pairs within a controlled number of matches.
-These win rates are then interpreted as Elo ratings, allowing us to create a comprehensive ranking of participating LLMs based on their relative performance.
-We validate the effectiveness of our approach, Varco Arena, against the current practice of LLM benchmarks, anchored comparison to reference, through two experiments: one is unbiased judge simulation and the other is running Varco Arena with LLM-as-a-judge on Arena-Hard-Auto benchmark dataset.
-These experiments confirm that our reference-free benchmarking approach outperforms current practices in terms of rank reliability, given the same number of comparisons.
-Notably, the simulation experiment, which emulates outcomes under an unbiased judge and controls for factors such as the number of participant LLMs, test prompts, and judge precision, demonstrates that the empirical superiority of our reference-free approach is not coincidental.
-Instead, it reveals the effectiveness of direct comparison between LLM pairs compared to the current practice of mediated comparison with references for ranking LLMs.
-To summarize, our key contributions are as follows:
+The need for LLM benchmarks extends beyond this, as the primary value of LLMs lies in the versatility of their generative behavior.
+Researchers have primarily relied on pairwise comparisons to evaluate and rank the LLMs based on human preference annotations.
+A representative example of this approach is Chatbot Arena Chiang et al.
+(2024), which computes Elo ratings based on a massive number of human votes.
+While Chatbot Arena benefited from reliable, open-ended prompts generated by a large user base, obtaining such a high volume of annotations from dedicated annotators remains challenging.
+Figure 1: VARCO Arena directly compares LLM response pairs in single-elimination tournament rather than comparing references.
+In terms of deciding whether a certain LLM is better or worse compared to the other one, we suggest direct head-to-head comparison is more intuitive and results in better separability.
+No. Comp. ($\downarrow$) Judge Eval. Type
+Chatbot Arena unknown human head-to-head
+Current Practice $n_{\text{model}}\cdot|X|$ LLM reference -based
+VARCO Arena $(n_{\text{model}}-1)\cdot|X|$ LLM head-to-head
+Table 1: Comparison between Current Practice and Varco Arena.
+$|X|$ and $n_{\text{model}}$ represents size of benchmark dataset, and number of candidate LLMs to rank respectively.
+Human annotators are considered much more costly than LLM judge counterpart.
+To address these challenges, several benchmark datasets for generation tasks have been developed and applied with an LLM judge Zheng et al.
+(2024) to quantify LLM capabilities.
+Using these benchmarks and their reference responses, LLMs are ranked via pairwise comparisons between their outputs and the reference responses provided by LLM judges.
+Representative examples include AlpacaEval Li et al.
+(2023), Arena-Hard-Auto Li et al.
+(2024), and MTBench Zheng et al.
+(2023).
+There are two major advantages to relying on reference responses: (1) The number of comparisons required for ranking scales linearly with the number of LLMs.
+(2) The reference response establishes a quality standard for evaluating LLM outputs.
+However, we argue that using reference responses to mediate comparisons between LLM outputs is suboptimal and that head-to-head comparisons yield more accurate assessments.
+In this context, we propose VARCO Arena as a novel evaluation framework that directly compares LLM responses while requiring fewer comparisons.
+Even with fewer comparisons and without relying on reference responses, VARCO Arena achieves a stronger correlation with the human-established Elo rankings from Chatbot Arena.
+VARCO Arena is designed to conduct a single-elimination tournament for each prompt across all participating LLM responses and then compute Elo ratings Elo and Sloan (1978) based on the outcomes of these matches.
+By organizing tournament-style matches for multiple LLMs on each prompt, we obtain relative win rates between all possible model pairs with fewer comparisons than those required by reference-based methods.
+These win rates are subsequently interpreted as Elo ratings, enabling us to generate a comprehensive ranking of candidate LLMs.
+We validate the effectiveness of VARCO Arena against current LLM benchmarking practices through two experiments.
+First, we conduct a simulation study (Section 4.2) to evaluate the reliability and robustness of our tournament approach under various controlled factors, including the number of participating LLMs, the number of test prompts, and the accuracy of the judge model.
+Second, we empirically demonstrate that running VARCO Arena with several judge LLMs on public benchmark (Section 4.3) prompts yields a stronger correlation with human-established Elo rankings from Chatbot Arena than does the reference-based approach.
+These experiments confirm that our benchmarking approach outperforms current practices in terms of ranking reliability while requiring the lesser number of comparisons (Table 1).
+Notably, the simulation results—which emulate outcomes under various conditions—underscore that the superiority of our tournament approach is not coincidental but rather reflects the inherent advantages of direct pairwise comparisons over mediated comparisons using reference responses.
+In this work, we make the following contributions:
   1. 1.
-    We introduce Varco Arena, a reference-free benchmarking method for LLMs that employs direct comparison of outputs through a single-elimination tournament structure across test prompts.
+    We introduce VARCO Arena, a novel reference-free benchmarking method for LLMs that leverages a single-elimination tournament structure to perform head-to-head comparisons of LLM responses across test prompts.
+This approach eliminates the dependency on static references and enhances benchmarking flexibility.
   2. 2.
-    We demonstrate, through conceptual validation and empirical results, that Varco Arena delivers more reliable rankings than current practices of comparing LLM outputs to reference texts, while requiring fewer comparisons.
+    Through extensive simulation studies and empirical evaluations, we demonstrate that VARCO Arena not only yields more reliable LLM rankings compared to traditional reference-based methods but also achieves this with significantly fewer comparisons, thereby reducing evaluation costs.
+  3. 3.
+    We release our demo and code to facilitate the further research and reproducibility; the code is available at [URL placeholder]
 2 Preliminaries: Quantifying Generation Ability
-Quantifying an LLM’s generation ability is crucial for ranking models, but it presents several challenges.
-The outcomes of comparisons between two LLMs are often probabilistic, influenced by various factors such as test prompts and the subjective nature of human preferences.
-A straightforward approach is to evaluate an LLM’s performance across diverse test prompts, which approximates its capability in real-world scenarios.
-Two popular measures for LLM generation ability are: win rate against reference text and elo rating.
-2.1 Win rate Against the Reference Text
-Several benchmarks quantify response generation quality, with AlpacaEval (Li et al.
-2023) and Arena-Hard (Li et al.
-2024) being representative examples.
-These benchmarks employ LLMs as judges for automated evaluation of generation capability.
-Given a prompt and two responses, the judge LLM is tasked with determining the superior response.
-These benchmarks include predefined sets of test prompts and reference outputs that serve as quality standards.
-The evaluation process involves comparing the LLM’s output against the reference to assess response quality.
-The LLM’s win rate across test prompts is used as a measure of its generation proficiency.
+Quantifying an LLM’s generation ability is crucial, but it presents several challenges.
+The outcomes of comparisons between two LLMs are often probabilistic, influenced by various factors such as the provided test prompts and the inherently subjective nature of human preferences.
+A straightforward approach is to evaluate an LLM’s performance across diverse test prompts, which approximates its real-world performance.
+Two popular measures for LLM generation ability are the win rate against reference responses and the Elo rating.
+2.1 Win Rate in Comparison to the Reference
+AlpacaEval Li et al.
+(2023) and Arena-Hard-Auto Li et al.
+(2024) are representative benchmark datasets that aim to quantify LLM response quality using reference responses.
+These benchmarks employ an LLM judge with a standardized prompt for automated evaluation of generation capability.
+Given the prompts, LLM responses, and reference responses, the judge LLM is tasked with determining whether the reference or the response is preferred.
+The LLM’s win rate across the test prompts in the benchmark is used as a measure of its generation proficiency.
 2.2 Elo Rating
@@
 While a superior team or LLM is not guaranteed to outperform its inferior counterpart in every instance, it tends to succeed more frequently.
-With Elo ratings, one can expect the relative chance of winning between a1 pair of players.
-The expected win rate between a pair of players $i$ and $j$ is computed as follows:
+With Elo ratings, one can expect the relative chance of winning between a pair of players.
+The expected win rate of the player $i$ against $j$ is computed as follows:
 […]
@@
 3 VARCO Arena
-Varco Arena represents our novel approach to reference-free benchmarking of LLMs.
-Rather than comparing LLM responses to references, we propose iterated tournaments over benchmark dataset to build leaderboards.
-This addresses concerns about effectiveness and scalability of using direct comparison over traditional benchmarking practice of relying on references.
-Section 3.1 and Algorithm 1 detail how Varco Arena runs iterated tournaments as we suggest.
-We then explain our rationale for using direct comparison of LLM responses over bridged comparing to references (Section 3.2), and argue why iterated tournaments can produce effective LLM leaderboards (Section 3.3).
-Algorithm 1 Running Iterated Tournaments over Benchmark for multiple LLMs
-1:Test prompts $X$, LLMs $M$, outputs $O_{ij}$ = $m_{j}(x_{i})$
+We propose VARCO Arena, a reference-free approach for benchmarking large language models (LLMs) via single-elimination tournaments.
+Instead of comparing model responses against fixed references, VARCO Arena directly compares outputs from different models, determining superiority through head-to-head matchups for each prompt in our benchmark dataset.
+Repeated tournaments across prompts yield reliable leaderboards that reflect the relative performance of each model.
+We begin by motivating our approach over current reference-based evaluations (Section 3.1).
+Next, we detail how VARCO Arena performs tournaments and results in Elo ratings (Section 3.2 and Algorithm 1).
+Finally, we discuss the reason why aggregating the results from tournaments promises reliable ranking (Section 3.3).
+3.1 Comparing to a Reference is not Always Helpful
+Although reference texts are a standard way to evaluate and rank large language models (LLMs), they introduce potential failure modes.
+Beyond the fact that a single reference might not capture every dimension of correctness, relying solely on a reference can lead to unreliable rankings of LLMs.
+Consider an ideal scenario with a judge capable of perfectly distinguishing the quality of any two outputs.
+If we choose to compare LLM responses directly to rank them using Elo ratings (Equation 1), all head-to-head comparisons are utilized.
+In contrast, reference-based evaluation for differentiating LLMs can exhibit failure modes, as shown in Equation 2.
+$\begin{array}[]{l}\text{M}_{1}(X_{i})\\ \quad\text{vs.}\,\rightarrow\\ \text{M}_{2}(X_{i})\end{array}\hfil\hskip 20.00003pt&\begin{cases}\text{M}_{1}(X_{i})>Y_{i}>\text{M}_{2}(X_{i})&\text{(helpful)}\\ \text{M}_{1}(X_{i})<Y_{i}<\text{M}_{2}(X_{i})&\text{(helpful)}\\ \text{M}_{1}(X_{i}),\ \text{M}_{2}(X_{i})>Y_{i}&\text{(unhelpful)}\\ \text{M}_{1}(X_{i}),\ \text{M}_{2}(X_{i})<Y_{i}&\text{(unhelpful)}\end{cases}\end{array}$
+When the reference output ($Y_{i}$) for a prompt ($X_{i}$) successfully disambiguates the pair of LLM responses $M_{1}(X_{i})$ and $M_{2}(X_{i})$ (as in the first and second cases), comparison to the reference is effective for benchmarking.
+Otherwise, these comparisons do not help differentiate LLM performance.
+Consequently, the reference-based approach provides less information for ranking when multiple responses are either both correct or both incorrect relative to the reference.
+3.2 Tournaments of LLMs over multiple prompts to Elo Ratings
+Algorithm 1 Tournaments of LLMs over prompts
+1:prompts $X=\{x_{1},~{}x_{2},~{}...,~{}x_{i}\}$, LLMs $M=\{m_{1},~{}m_{2},...,~{}m_{j}\}$, outputs $O_{i,j}$ = $m_{j}(x_{i})$
 2:Ranked LLMs with Elo ratings
 3:function Match($m_{1},m_{2},x$)
-4:     if $\text{IsBetter}(O_{x,m_{1}},O_{x,m_{2}})$ then
-5:         return $m_{1}$
-6:     else
-7:         return $m_{2}$
-8:     end $\triangleright$ omitted byes for conciseness
-9:end function
-10:function SingleElim(players, $x$, results)
-11:     if ||||players|=2|=2| = 2 then
-12:         $m_{\text{win}}$leftarrow Match(players[0], players[1], $x$)
-13:         results.append($m_{\text{win}}$)
-14:         return $m_{\text{win}}$
-15:     end if
+4:     return $m_{1}$ if $\texttt{IsBetter}(O_{x,1},O_{x,2})$
+5:     else $m_{2}$
+6:end function
+7:function SingleElim($M$, $x$, res)
+8:     if $|M|=2$ then
+9:         res.append(Match($M$[0], $M$[1], $x$))
+10:         return res[-1]
+11:     end if
+12:     mid $\leftarrow\lfloor|M|/2\rfloor$
+13:     left $\leftarrow$ SingleElim($M$[:mid], $x$, res)
+14:     right $\leftarrow$ SingleElim($M$[mid:], $x$, res)
+15:     return SingleElim(left + right, $x$, res)
+16:end function
+17:function Tournaments2Ranks($X,M$)
 […]
-17:     l_bracket $\leftarrow$ SingleElim(players[:mid], $x$, results)
-18:     r_bracket $\leftarrow$ SingleElim(players[mid:], $x$, results)
-19:     return SingleElim(l_bracket + r_bracket, $x$, results)
-20:end function
-21:function Tournament2Leaderboard($X,M$)
-22:     results $\leftarrow$ []
-23:     for $x_{i}\in X$ do
-24:         SingleElim(Shuffled($M$), $x_{i}$, results)
-25:     end for
-26:     return ComputeElo(results)
-27:end function
-3.1 Iterating Tournaments over Benchmark Dataset to Elo ratings
-Figure 1 and Algorithm 1 illustrate the methodology of Varco Arena for benchmarking Large Language Models (LLMs) using a tournament approach.
-In this context, $|X|$ denotes the cardinality of the prompt set in the benchmark dataset (i.e.
-number of prompts in the dataset).
-As described running Varco Arena invokes tournament of participant LLMs for each prompt in benchmark dataset, which we will refer to as iterated tournaments.
-The application of tournament structures to LLM benchmarking presents both opportunities and challenges.
-While a single tournament effectively identifies a champion, it does not provide the granularity required for a comprehensive ranking system; it only declares a champion, leaving the precise ordering of other participants ambiguous.
-To retain the efficient nature of tournaments but to accomplish a fine-grained ranking of LLMs, what modifications should be applied to?
-To mitigate the aforementioned ranking resolution issue, we propose conducting independent tournaments for each benchmark prompt, with randomized initial match-ups.
-This approach, which we term iterated tournaments, offers several advantages:
+19:     for $x_{i}\in X$ do
+20:         SingleElim(Shuffled($M$), $x_{i}$, res)
+21:     end for
+22:     return ComputeElo(res)
+23:end function
+Figure 1 and Algorithm 1 illustrate how VARCO Arena benchmarks LLMs via a tournament approach.
+Here, $|X|$ denotes the number of prompts in the benchmark dataset.
+Each execution of VARCO Arena runs a tournament among participant LLMs for every prompt in the dataset.
+The use of tournament structures for LLM benchmarking offers both benefits and challenges.
+A major advantage of a single-elimination tournament is efficiency.
+As shown in Table 1, the number of matches scales linearly with the number of participants and even lower compared to using references.
+However, single elimination tournament only identifies a champion, leaving the relative ordering of other participants unclear.
+To retain tournament’s efficiency while obtaining a fine-grained ranking, we propose aggregating tournament results over multiple prompts with randomized initial match-ups for each prompt.
+Performing multiple tournaments with random initialization offers several benefits:
   1. 1.
-    It facilitates the resolution of ties among non-champion participants from previous tournaments.
+    It resolves ties among non-champion participants from previous tournaments.
   2. 2.
-    It reduces the impact of unfavorable match-ups in any single tournament.
+    It mitigates the impact of unfavorable match-ups in any single tournament.
   3. 3.
-    While not providing exhaustive pairwise comparisons, the accumulated matches between model pairs enable more precise estimation of relative win rates and Elo ratings, culminating in a well-aligned overall ranking.
-In terms of total number of matches, our iterated tournaments approach requires $|X|*(n_{\text{models}}-1)$ which is always $|X|$ numbers of comparisons saved compared to anchored comparison to references.
-A favorable trait of conducting tournaments for each prompt is that every participating LLM engages in at least one competition per prompt.
-The number of matches an LLM faces per test prompt varies from 1 to $\lceil\text{log}_{2}(n_{\text{models}})\rceil$, based on its performance.
-Allocating more matches for better LLMs will help precise separation between higher rated models.
-We further elaborate and argue why aggregating multiple tournaments should reach a good sorting of participant LLMs in Section 3.3.
-Additionally, we provide prospect match numbers LLMs would confront within, offering a comprehensive view of the method’s efficiency and effectiveness.
-3.2 Comparing to Reference is not Always Helpful
-Here, we inspect into failure modes of relying on references for ranking LLMs, which further elaborates the previously mentioned concerns on current practices for LLM benchmarking.
-Consider a highly accurate judge capable of definitively comparing two outputs of different quality without errors or ties.
-In this scenario, all matches involving participant LLMs contribute to their relative win rates and, ultimately, to their final Elo ratings (Section 3.3 and Equation 2).
-Consequently, all matches involving the LLM of interest contribute to Elo rating and ranking of the LLM.
-When evaluating LLM performance relative to reference texts, even with the same precise judge, the comparison doesn’t always provide a reliable measure of relative win rates.
-Suppose we have only two participants of interest $\text{M}_{1}$ and $\text{M}_{2}$ and an instruction $X_{i}$ to prompt these LLMs.
-There might be two corresponding responses $\text{M}_{1}(X_{i})$, $\text{M}_{2}(X_{i})$ and also a reference text $Y_{i}$.
-Equation 2 illustrates four possible scenarios.
-In two of these cases, the match outcome is useful for comparing two LLMs, specifically when the reference text loses to one model and wins against the other, working as a useful anchor for comparison.
-In the other two cases, comparing to $Y_{i}$ fails to differentiate between $\text{M}_{1}(X_{i})$ and $\text{M}_{2}(X_{i})$.
-$\begin{array}[]{l}\text{M}_{1}(X_{i})\\ ~{}~{}~{}~{}~{}\text{vs.}~{}~{}~{}~{}~{}\rightarrow\\ \text{M}_{2}(X_{i})\end{array}\begin{cases}\text{M}_{1}(X_{i})>Y_{i}>\text{M}_{2}(X_{i})&\\ ~{}~{}~{}~{}~{}\text{or}&\text{useful}\\ \text{M}_{1}(X_{i})<Y_{i}<\text{M}_{2}(X_{i})&\\ &\\ \text{M}_{1}(X_{i}),~{}\text{M}_{2}(X_{i})>Y_{i}&\\ ~{}~{}~{}~{}~{}\text{or}&\text{useless}\\ \text{M}_{1}(X_{i}),~{}\text{M}_{2}(X_{i})<Y_{i}&\\ \end{cases}$ (2)
-This analysis reveals a potential weakness in benchmarks that rely on anchored comparisons to reference texts.
-If the reference texts are not carefully calibrated in terms of difficulty, the ranking of LLMs could be determined by only a small portion of the test prompts.
-Specifically, if reference texts are consistently superior or inferior to most model outputs, a large number of comparisons become uninformative.
-This issue is mitigated when directly comparing model outputs without a reference text, suggesting that direct comparisons may provide a more robust method for ranking LLMs.
-3.3 Iterated Tournaments as a Rough Approximation of Merge Sort
-In this section, we argue that iterated tournaments will reach to a good sorting of participants LLMs comparing the similarities in merge sort algorithm and iterated tournament regarding the presumptions of Elo model for LLM capability.
-Our approach is predicated on a key assumption: although the prompts in the benchmark dataset are diverse, they can be considered approximately equivalent in their capacity to differentiate LLM performance.
-This assumption of ‘indistinguishable prompts’ aligns with the Elo model of LLM capability, which posits that prompt-specific variations should not significantly affect the expected outcome likelihood of a single match.
-[^1] Given the simplicity of the Elo model’s assumptions regarding LLM capability, we posit that ranking LLMs with Elo ratings for a mixture of disparate capabilities may be inappropriate, as a higher Elo rate does not guarantee uniform mastery across multiple capabilities.
-Building on this assumption, we propose that conducting tournaments across the benchmark prompts can lead to an accurate ranking of LLMs.
-We suggest that single-elimination tournaments function analogously to merge sort, with a crucial distinction: they perform only the initial comparison at each step of a merge operation and discard the losing side (i.e.
-only the winner proceeds to the next rounds).
-This makes a single tournament perform $n-1$ comparisons in total.
-This number of comparisons is less than $\mathcal{O}(n\text{log}_{2}n)$ which most of the efficient sorting algorithms requires.
-However, regarding the assumption of ‘indistinguishable prompt’ we mentioned above, number of tournament matches from the other dataset prompt would help filling the lack of comparion matches.
-By iterating over the benchmark dataset with randomized initial match-ups, these tournament matches can eventually approximate the complete set of comparisons required for merge sort, given a sufficient number of benchmark prompts ($|X|$).
-Our approach provides $\mathcal{O}(|X|$/$\text{log}_{2}n_{\text{model}})$ times more comparisons [^2] Where $\mathcal{O}(|X|$/$\text{log}_{2}n_{\text{model}})=\frac{|X|(n_{\text{model}}-1)}{\mathcal{O}(n_{\text{model}}\text{log}_{2}n_{\text{model}})}$ representing the ratio between matches in iterated tournaments and matches needed for sorting LLMs with an ideal judge.
-than the minimum required for merge sorting $n_{\text{model}}$ LLMs.
-In practice, number of matches to figure which model is better from a pair is much larger since the judge model will not be ideal and also performance of LLMs will be stochastic regarding the prompt it was asked to answer.
-Nevertheless, the large number of prompts ($|X|$), typically ranging from hundreds to thousands, lead us to the worth of empirical validation of the idea, that iterated tournaments will accurately sort LLMs with confidence.
-This foundational argument underpins our unbiased judge simulation experiment to test the potential of tournament approach (Varco Arena).
+    Aggregating match results allows for precise win rate estimation via Elo ratings, resulting in a well-aligned overall ranking.
+  4. 4.
+    More matches are allocated to high-performing participants while ensuring every participant is evaluated at least once per prompt.
+In Section 3.3, we further explain how aggregating multiple tournaments could yield an reliable ranking of LLMs.
+We also provide an analysis of the number of matches each LLM faces, offering a comprehensive view of the method’s efficiency and effectiveness.
+3.3 Why Aggregating Multiple Tournaments Yields Reliable Ranks?
+Our approach aggregates match outcomes from multiple tournaments to approximate the complete set of pairwise comparisons—akin to the comparisons made in merge sort.
+In a single-elimination tournament, every participant advances based solely on match outcomes, a process that mirrors the merging steps in merge sort.
+Notably, a single-elimination tournament executes only the comparisons strictly necessary for determining the winner, omitting many comparisons that merge sort would perform further.
+We posit that the missing pairwise match-ups in any one tournament can be recovered by aggregating tournaments conducted over different prompts.
+This hypothesis relies on the assumption—central to our use of the Elo model—that match outcomes are independent of the prompt.
+Consequently, matches across different prompts are considered equivalent.
+Considering only the initial comparisons, which are randomly sampled, the aggregate number of comparisons is $|X|\cdot n_{\text{model}}/2$.
+Since $n_{\text{model}}$ is typically on the order of tens and $|X|$ comprises at least hundreds, this number exceeds the total possible match-ups, $\binom{n_{\text{model}}}{2}$ [^1] Note that merge sort produces a set of match-ups with no duplicates..
+Furthermore, as shown in Table 1, the remaining matches—totaling $|X|\cdot(n_{\text{model}}-1)$—either contribute additional merge sort structures or enhance the accuracy of estimating relative win rates among LLM participants.
+Moreover, considering that each unique pair meets in at least $|X|/(2(n_{\text{model}}-1))$ matches [^2] A model participates in between $|X|$ and $\lceil\log_{2}n_{\text{model}}\rceil$ matches per tournament.
+Dividing this by the number of possible match-ups per model, $2(n_{\text{model}}-1)$, yields the expression above.
+across the benchmark, this frequency is sufficient for a fair estimation of their relative win rates.
+In summary, aggregating tournaments not only reconstructs the full set of pairwise comparisons for a merge sort but also ensures that each pair of models faces one another often enough to yield accurate win rate estimations, leading to reliable Elo ratings.
+Based on these considerations, we propose that conducting tournaments over a benchmark prompts will yield a reliable ranking of LLMs.
 4 Experiments
-We run two experiments with different settings for comparing tournament and anchored comparison approaches.
-The first experiment (Section 4.2) simulates an unbiased judge to control various factors that affect LLM benchmarking.
-This simulation tests our foundational propositions for Varco Arena design (mentioned in Section 3.2 and 3.3) under a more controlled, simplified environment.
-It eliminates noisy factors such as potential biases of LLM judges (Park et al.
-2024).
-In subsequent empirical experiments (Section 4.3), we use gpt-4o[-mini] as a judge, to validate the effectiveness of our tournament approach against current LLM benchmarking practices.
+We run two experiments with different settings for comparing VARCO Arena and current practice of reference-based benchmarking.
+In the first experiment (Section 4.2) we conduct a simulation study to control various factors that affect LLM benchmarking.
+This simulation tests our foundational propositions for VARCO Arena design (mentioned in Section 3.1 and 3.3) under a more controlled, simplified environment immune to noisy factors such as potential biases of LLM judges Park et al.
+(2024).
+The other is empirical experiments (Section 4.3).
+We use gpt-4o and -mini as well as other several popular models such as Claude3.5, Qwen2.5, Llama3.1, and Gemma2 to validate the effectiveness of our tournament approach against current LLM benchmarking practices.
 By presenting both simulation and empirical results, we aim to demonstrate the effectiveness of our tournament approach.
-We first describe the common experimental settings before getting into specific details in the following subsections.
+In Section 4.1, we first describe the common experimental settings before getting into specific details of each experiments in the following subsections.
 4.1 Chatbot Arena Leaderboard Ratings as Ground-Truth LLM Rankings
 We compare the results of each benchmarking approach against the rankings from the Chatbot Arena leaderboard.
-Chatbot Arena is currently considered one of the most reliable leaderboards due to its large number of human preference annotations.
-In this platform, users are presented with responses from a pair of LLMs to the same query and asked to choose which response they prefer.
-Given the substantial number of votes and the wide range of submitted prompts used for model comparison, the resulting ranks from Chatbot Arena are likely accurate enough to serve as ground-truth rankings.
-4.2 Experiment 1: Simulating an Unbiased Judge
-We designed a simple stochastic modeling experiment to simulate LLM matches following the Elo preference model.
-Our model assumes that the primary factor governing the outcome of each match is the Elo rating gap ($\Delta_{\text{ij}}$), as shown in Equation 3.
-We also introduced a secondary factor, judge precision ($P_{\text{judge}}$), which represents the likelihood that a judge correctly declares the higher-rated LLM as the winner.
-The observed outcome of a single match is sampled according to the likelihood of model $i$ winning against model $j$ ($P_{\text{predict}}(i>j)$).
-This probability is the product of the judge precision and the estimated relative win rate expected by ground-truth Elo ratings ($P_{\text{gt}}(i>j)$).
-This simplified simulation model aligns with the fundamental principle of the Elo system: the outcome of a match between a pair of models is primarily determined by their skill gap ($\Delta_{\text{ij}}$).
-This gap is derived from past wins and losses, effectively condensing all sophisticated factors into a single mean value.
-$P_{\text{predict}}(i>j)=P_{\text{judge}}*P_{\text{gt}}(i>j)=P_{\text{judge}}*\frac{1}{1+10^{\Delta_{\text{ij}}/400}}$ (3)
+Chatbot Arena is widely regarded as one of the most reliable leaderboards due to its extensive collection of human preference annotations.
+Given the large number of votes and the diverse set of prompts used for model comparisons, the resulting rankings are considered sufficiently accurate to serve as ground truth.
+4.2 Experiment 1: Simulation Study
+We designed a simple simulation to emulate a probabilistic model of LLM matches, adhering to the Elo rating system in a controlled environment.
+In line with the Elo model’s assumptions, our judge is configured to follow Equation 3 exactly.
+The judge stochastically determines the winner of each LLM match solely based on the Elo rating difference ($\Delta_{\text{ij}}$) and the judge’s accuracy ($P_{\text{judge}}$).
+As described in Equation 3, the outcome of a single match is sampled according to $P_{\text{predict}}(i>j)$, which is computed as the product of the judge’s accuracy and the likelihood of model $i$ beating model $j$ based on the Elo gap ($P_{gt}$).
+$\begin{split}P_{\text{predict}}(i>j)&=P_{\text{judge}}\times P_{\text{gt}}(i>j)\\ &=P_{\text{judge}}\times\frac{1}{1+10^{\Delta_{\text{ij}}/400}}\end{split}$ (3)
 The details of our simulation settings are as follows:
-Ground-truth Elo ratings: We extracted Elo ratings from the English category of Chatbot Arena as of June 23, which includes one million user-submitted prompts and judgments (approximately 60% of the total submissions to the platform).
-While any set of Elo ratings could be valid for this simulation, we chose to use real values computed from human preferences.
-Judge Precision: In real-world, judge precision would be an adaptive value that varies based on the prompt-response pair being judged or the method of prompting the LLM-as-a-judge.
-In our experiments, we controlled judge precision, ranging from 0.6 to 0.9.
-Number of participant LLMs ($n_{\text{model}}$) and benchmark dataset size ($|X|$): We varied these factors to assess the robustness of tournament and anchored comparison approaches under diverse conditions.
-A larger number of participants may require more matches to achieve precise results.
-Similarly, the benchmarking environment may be data-poor or data-rich.
-We emulated the varying dataset size by sampling different numbers of matches from $P_{\text{predict}}$.
-The simulation process follows these steps:
+Ground-truth Elo ratings (initial parameter of the simulation, and at the same time, the gold ranking to reproduce): We extracted Elo ratings from the English category of Chatbot Arena as of June 23.
+This Elo ratings are the estimates from massive user-submitted judgments (approximately 60% of the total submissions to the platform).
+For the simulation, any set of Elo ratings could be used, but we opted for real values computed from human preferences.
+Judge Accuracy ($P_{\text{judge}}$): In practice, a judge’s accuracy is an adaptive value that depends on factors such as the specific prompt-response pair and the manner in which LLM judges are prompted.
+In this simulation study, we control this parameter, varying it from 0.6 to 0.9.
+Number of Participant LLMs ($n_{\text{model}}$) and Benchmark Dataset Size ($|X|$): We varied these factors to assess the robustness of both the tournament and reference-based approaches under diverse conditions.
+This allowed us to explore how reliability changes with the number of participants in both data-poor and data-rich environments.
+Simulation Procedure:
   1. 1.
-    Select participants from the top of the Chatbot Arena leaderboard, avoiding ties (e.g., selecting only one model from a tied rank).
+    Select the participant LLMs and obtain their Elo ratings for the simulation.
   2. 2.
-    Compute expected relative win rates ($P_{\text{gt}}$, Equation 3) from the participants’ Elo ratings.
+    Compute the expected relative win rates ($P_{\text{gt}}$, see Equation 3) from the participants’ Elo ratings.
   3. 3.
-    Sample match outcomes across LLM pairs for tournaments or anchored comparisons.
-An unbiased judge with precision $P_{\text{judge}}$ (Equation 3) determines the winner according to Equation 3, considering only the Elo gap $\Delta_{\text{ij}}$.
+    Sample match outcomes for each LLM pair according to the benchmarking approach.
+The winner of each match is sampled from $P_{\text{pred}}$ (Equation 3), which depends solely on the Elo gap ($\Delta_{\text{ij}}$) and the judge’s accuracy ($P_{\text{judge}}$).
   4. 4.
-    Repeat step 3 for the designated benchmark dataset size $|X|$.
+    Repeat step 3 for the designated number of test prompts ($|X|$).
   5. 5.
@@
     1. (a)
-      For anchored comparison, use the win rate against the reference LLM (we choose gpt-4-1106-preview which records 1233 in Elo rate).
+      For the reference-based approach (current practice), use the win rate against the reference model (gpt-4-1106-preview, which has an Elo rating of 1233).
     2. (b)
-      For tournaments, compute Elo ratings from the recorded match outcomes from every tournaments performed.
+      For the tournament approach (VARCO Arena), compute Elo ratings from all match outcomes.
   6. 6.
-    Rank models based on these scores.
+    Rank the models based on these scores.
   7. 7.
-    Compute the Spearman correlation between the simulated rankings (6) and the ground-truth rankings (1).
-We perform 50 trials per simulation configuration to mitigate randomness from tournament brackets and sampling.
-4.3 Experiment 2: Running Varco Arena using gpt-4o[-mini] as Judge
-We conducted an empricial experiemnt of Varco Arena to validate our proposed method.
-We conducted tests for both tournament approach and anchored comparison approach using 20 top-ranked models from the Chatbot Arena leaderboard.
-This experiment differs from our modeling simulation in several key aspects.
-We elaborate on the details of the experiments in order of how we prompted the LLM judges and our practice to managing API costs while conducting a large number of experiments.
-Dataset for Testing Benchmarking Approaches
-In order to test the benchmarking approaches, tournament and anchored comparison, we need two sorts of data.
-One is a prompt to be asked for LLMs and the other is a reponse by each LLM given the prompt.
-For the benchmark dataset, we choose Arena-Hard-Auto (Li et al.
-2024).
-Arena-Hard-Auto benchmark prompts are carefully assorted from Chatbot Arena which are user-typed queries to LLMs.
-The total number of prompts in Arena-Hard-Auto is 500, each 2 instances for 250 subtopics.
-Choosing Arena-Hard-Auto as our benchmark dataset is natural since it minimizes the disparity between our benchmarking environment and the ground truth environment where human annotators actually judged LLMs.
-As suggested in Arena-Hard-Auto, we use the responses of gpt-4-0314 as reference outputs.
-For responses of the LLMs to be ranked, we use the reserved outputs of the top 21 models from Huggingface space of the Arena-Hard-Auto Browser.
-[^3] Extracted from Jul 6 commit (fd42026)
-Participant LLMs
-As a list of models to be ranked, we choose 20 LLMs from the top of ChatBot Arena leaderboard in Hard prompts category which is considered to synchronize the most to Arena-Hard-Auto.
-To avoid egocentric bias, we exclude judge models from the participant LLMs.
-We use gpt-4o-2024-05-13 and *4o-mini as our judge, gpt-4-0314 as a reference (anchor) model to perform anchored comparisons on.
-You can see the evaluation prompt on Listing 1 at Appendix.
-LLM Judges
-We deploy gpt-4o and gpt-4o-mini as judges.
-To prompt the pairwise comparison of the responses from the participant LLMs, we refer to (Zeng et al.
-2024), equally for both tournament and anchored comparison approaches.
-To avoid position bias (Wu and Aji 2023), we alternate the position of the model responses for prompting evaluation.
+    Compute the Spearman correlation between the simulated rankings (from step 6) and the ground-truth rankings (from step 1).
+We perform 50 trials for each simulation configuration to mitigate randomness from tournament brackets and sampling.
+4.3 Experiment 2: Assessing VARCO Arena Using Various LLM Judges
+To empirically validate our proposal, we evaluated the reliability of both VARCO Arena and reference-based approach over the top 20 models from the Chatbot Arena leaderboards.
+This experiment employs actual prompt inputs and LLM outputs, distinguishing it from the earlier simulation study.
+4.3.1 Dataset: Test Prompts and LLM Responses Used
+Testing the benchmarking approaches requires: (1) test prompts and (2) the corresponding responses from LLMs.
+For the benchmark dataset, we selected Arena-Hard-Auto Li et al.
+(2024).
+The prompts in Arena-Hard-Auto were carefully curated from Chatbot Arena user queries.
+This dataset consists of 500 prompts—two instances for each of 250 subtopics.
+Although AlpacaEval Li et al.
+(2023), which comprises 800 prompt-reference pairs, could serve as a viable testbed, we opted for Arena-Hard-Auto because its design aligns more closely with Chatbot Arena.
+Arena-Hard-Auto uses responses from gpt-4-0314 as the reference outputs.
+For ranking, we utilized the reserved outputs of the top 21 models from the Arena-Hard-Auto Browser.
+[^3] Extracted from the 2024 Jul 6 commit (fd42026).
+4.3.2 Participant LLMs
+For ranking, we selected 20 LLMs from the top of the ChatBot Arena leaderboard in the hard prompts category, as these models most closely align with Arena-Hard-Auto.
+4.3.3 LLM Judges
+We used several aligned LLMs as judges for testing both benchmarking approaches.
+LLMs of our choice are gpt-4o family of models OpenAI et al.
+(2024), Claude3.5, and a selection of open-weight models: Qwen2.5 Qwen et al.
+(2025), Llama3.1 Grattafiori et al.
+(2024), and Gemma2 Team et al.
+(2024).
+For pairwise comparisons of responses, we employed the judging prompt suggested in LLMBar Zeng et al.
+(2024) (See Appendix A.6.2).
+The same judge prompt was applied consistently across both the tournament and reference-based approaches.
+To mitigate position bias Wu and Aji (2023), the order of model responses was alternated during evaluation.
 Further details on the LLM-as-a-judge configuration are provided in Appendix A.6.
-While using gpt-4o-2024-05-13 as a judge provides a realistic setting for LLM benchmarking, we also explore gpt-4o-mini-2024-07-18 to accommodate users with budget constraints.
-Assuring Statistical Significance of the Results within Budget
+The two experimental settings are summarized as follows:
+Experiment 1 (Simulation Study): This experiment uses the ground truth Elo ratings of the models to initialize the simulation.
+We vary control parameters for the benchmarking approaches—including the judge’s accuracy ($P_{\text{judge}}$), the number of test prompts used ($|X|$), and the number of participant LLMs ($n_{\text{model}}$)—to determine which benchmarking approach more accurately reproduces the participants’ ranking.
+For each configuration, we conduct 50 trials of experiments.
+Experiment 2 (Empirical Runs): This experiment assesses the two benchmarking approaches using empirical runs with various LLM judges.
+We select the top 20 LLMs from Chatbot Arena and used their reserved outputs on Arena-Hard-Auto test prompts.
+For both the tournament and reference-based approaches, we employ the Spearman correlation coefficient to measure how well the results align with the ground truth leaderboard rankings.
+In our empirical study, we conduct 500 trials for each experimental setting.
+Figure 2: Simulation results comparing the tournament and reference-based approaches.
+The tournament method consistently outperforms the reference-based approach in Spearman correlation across various control variables: the number of participant LLMs ($n_{\text{models}}$), the number of benchmark prompts ($|X|$), and judge precision ($P_{\text{judge}}$).
+5 Results and Discussion
+We assess the reliability and robustness of VARCO Arena as a means for LLM benchmarking, comparing it against the current reference-based approach.
+Our results from both simulation study and empirical runs indicate that the tournament approach of VARCO Arena yields rankings that align more closely with the ground-truth Elo leaderboards.
+We present our findings using whisker plots and tables in the following sections.
+5.1 Experiment 1: Simulation Study Results
+Figure 2 illustrates noticeable differences in Spearman correlation, indicating that the tournament approach is more reliable than the reference-based method.
+The consistent performance gap across various conditions—namely, the number of participants, the number of test prompts, and judge accuracy ($n_{\text{model}}$, $|X|$, and $P_{\text{judge}}$)—demonstrates the robustness of the tournament approach.
+Although the simulation simplifies real-world complexity, a similar performance gap was observed in the empirical findings (Experiment 2, Figure 3).
+This consistency suggests that the robust performance of VARCO Arena is not coincidental or limited to a specific empirical setting of ours.
+5.2 Experiment 2: Empirical Results
+Figure 3: Results of VARCO Arena (tournament) and reference-based approach with gpt-4o (left) and gpt-4o-mini (right) judge.
+VARCO Arena constantly records higher Spearman correlation coherent with the Experiment 1 result (Figure 2).
+Results summary is on Table 2.
+As hinted in the previous section, the empirical results in Figure 3 show that VARCO Arena consistently outperforms the reference-based approach.
+Although the performance gaps are less pronounced than in the simulation, the same trend persists.
+In Table 2, we report the median values for VARCO Arena and the reference-based approach using the gpt-4o family of judges while varying the number of test prompts ($|X|$).
+These results consistently demonstrate that VARCO Arena outperforms the reference-based method.
+Note that VARCO Arena shows similar or superior reliability even in extreme data-poor benchmark condition ($|X|=50$).
+Table 3 presents the outcomes when using other LLMs as judges, with a fixed number of prompts ($|X|=500$).
+The results for Claude3.5-sonnet, Llama3.1-8b, and Qwen2.5-7b follow a similar trend.
+However, smaller models (Gemma2-2b and Qwen2.5-0.5b) appears to be less reliable for benchmarking.
+Hence, we recommend using evaluation-specialized judge LLMs or, at least, generative judge models with around 7B parameters regardless of using VARCO Arena or considering reference-based approach.
+Spearman corr. ($\uparrow$) $|X|$ = 50 100 250 475 500
+comp. to ref. (4o) 0.895 0.935 0.963 0.966 0.964
+tournament (4o) 0.905 0.940 0.960 0.970 0.970
+comp. to ref. (4o-mini) 0.895 0.908 0.917 0.916 0.912
+tournament (4o-mini) 0.901 0.919 0.931 0.933 0.933
+Table 2: Spearman correlation ($\uparrow$) varying over size of the benchmark set ($|X|$) for each benchmarking approach. Comp. to ref. refers to reference-based approach.
+$|X|$ = 500 claude3.5 sonnet llama3.1 8b-it qwen2.5 7b-it qwen2.5 0.5b-it gemma2 2b-it
+comp. to ref. 0.924 0.820 0.756 0.089 0.592
+tournament 0.930 0.850 0.811 -0.124 0.552
+Table 3: Spearman correlation ($\uparrow$) result using other LLMs as a judge. Comp. to ref. refers to reference-based approach.
+5.3 Incorporating a New LLM into an Existing Leaderboard
+While our main focus has been on ranking multiple LLMs at once, it is also useful to consider the common scenario of adding a single new model to an existing leaderboard, which is also frequent use-case for leaderboards.
+We explored two approaches: (1) a binary search-like placement method, and (2) using the top-performing model response as a reference.
+Our findings indicate that the latter approach is more reliable (Table 4).
+Further details and discussions are provided in Appendix A.4.
+$\bar{|\Delta_{\text{rank}}|}$ ($\downarrow$) gt=1-6 7-13 14-19 (20) total avg.
+binary search (4o) 0.92 1.84 2.13 1.72
+comp. to 1st (4o) 1.98 1.55 1.57 1.39
+binary search (4o-mini) 1.27 1.82 1.21 1.5
+comp. to 1st (4o-mini) 1.00 1.43 1.43 1.37
+Table 4: Comparison of the binary search method versus using the top-performing model’s response as a reference (comp.
+to 1st) for inserting a new LLM into the leaderboard.
+We report the mean rank deviation ($\bar{|\Delta_{\text{rank}}|}$) from the ground-truth leaderboard as an additional error metric.
+For further details, see Algorithm 2 in Appendix.
+6 Related Works
+6.1 Elo-based LLM Benchmarking
+Recent studies have leveraged Elo ratings derived from human preferences to benchmark LLMs Boubdir et al.
+(2023).
+For instance, Chatbot Arena Chiang et al.
+(2024) employs Elo as an evaluation metric, while RAGElo Rackauckas et al.
+(2024) uses it as a ranking metric.
+These implementations highlight the benefits of applying Elo ratings for open-ended text quality evaluation.
+6.2 Reference-free Evaluation
+Reference-free evaluation has emerged as a promising alternative to static reference-based methods from the era of Natural Language Generation.
+Advances in LLM capabilities have enabled models to assess open-ended responses effectively Jauhiainen and Guerra (2024).
+When reference quality is poor, reference-free metrics like XComet Guerreiro et al.
+(2023) works as a more reliable alternatives.
+7 Conclusion
+We introduced VARCO Arena, a reference-free LLM benchmarking approach that employs a tournament-style framework with direct pairwise comparisons to evaluate the generative capabilities of LLMs.
+VARCO Arena offers a cost-effective, scalable, and adaptable solution for benchmarking response quality.
+Our simulation study and empirical experiments demonstrate that VARCO Arena consistently achieves higher rank reliability compared to current reference-based methods, as evidenced by stronger correlations with human-established Elo ratings.
+Given its robust performance and flexibility, we believe VARCO Arena can serve as a reliable automated tool for model selection and ranking across diverse and evolving use cases.
+Future work will explore broader applications, such as benchmarking LLMs in multi-modal settings (e.g., those incorporating visual or audio inputs and outputs).
+Limitations
+Although we tested robustness of the tournaments performed by VARCO Arena, it still has added factor for randomness such as match bracket initialization which does not applies to reference-based method.
+Rooms for improvement exist for more informative match-making algorithm that would achieve better ranking than single-elimination tournaments within same or less number of matches.
+Appendix A Appendix
+A.1 Machine Requirements for Experiments
+Except the part we inferenced open-weight models such as Llama, Qwen and Gemma, our experiments are mostly do not require GPU usage.
+Inference are done on one A100 GPU, but T4 would be enough for reproducing our experiments.
+Otherwise, our experiments require querying API and post-processing those with CPU.
+Experiments could be run on personal desktops.
+The lowest specification of the machine we deployed had i5-8400 CPU, 16 GiB RAM.
+A.2 Assuring Statistical Significance of the Results within Budget for proprietary models
 To ensure a statistically significant number of trials for each experiment while staying within budget, we utilize OpenAI’s Batch API to prepare full-grid match outcomes (i.e., all-play-all matches for every prompt) in a cache file, allowing us to reuse these outcomes.
@@
 When experimenting with a subset of the Arena-Hard-Auto benchmark ($|X|<500$), we sample a stratified subset of the benchmark dataset for each new trial.
-To summarize the differences in settings: Experiment 1 (simulating unbiased judge) uses ground truth Elo ratings from Chatbot Arena (English category) and control parameters of our choices (e.g., $P_{\text{judge}}$, $|X|$, and $n_{\text{model}}$) to predict LLM rankings, which are then compared to the ground-truth rankings.
-Experiment 2 (running Varco Arena with LLM judges) benchmarks the top 20 models from Chatbot Arena (Hard prompts category) using the Arena-Hard-Auto benchmark dataset.
-For both tournament and anchored comparison approaches, we employ Spearman correlation as a correlation measurement to see how well the benchmarking approaches aligns to the ground-truth leaderboard.
-5 Results and Discussion
-We report on the reliability of Varco Arena for simultaneously ranking multiple LLMs.
-Both simulated and empirical results demonstrate that the tournament approach of Varco Arena produces ranks more closely aligned with ground-truth Elo ranks.
-5.1 Experiment 1: Unbiased Judge Simulation
-Figure 2: Simulated results comparing tournament and anchored comparison approaches with an unbiased judge.
-The tournament approach consistently demonstrates superiority in terms of Spearman correlation across various control variables: number of participant LLMs ($n_{\text{models}}$), number of prompts in the benchmark set ($|X|$), and judge precision ($P_{\text{judge}}$).
-Figure 2 reveals distinct gaps in Spearman correlation, indicating that the tournament approach demonstrates greater reliability compared to the anchored comparison approach.
-These significant gaps persist across various controlled parameters, including the number of participating LLMs ($n_{\text{models}}$), prompts in the benchmark set ($|X|$), and judge precision ($P_{\text{judge}}$).
-While the simulation necessarily abstracts real-world complexity, the consistent performance patterns observed under systematically varied conditions align with the empirical findings from Experiment 2, suggesting potential robustness of the tournament approach beyond the specific conditions tested empirically.
-5.2 Experiment 2: Running Varco Arena with gpt-4o[-mini] Judge
-Figure 3: Tournament and anchored comparison approach run with gpt-4o (left) and gpt-4o-mini (right) judge.
-Tournament approach constantly records higher Spearman correlation across the varying number of benchmark prompts used.
-Results summary is on Table 1.
-Our empirical experiments corroborate the trends observed in the simulation.
-As illustrated in Figure 3 and Table 1, while the gaps are less pronounced compared to the simulation results (Figure 2), the Varco Arena results obtained through the tournament approach consistently demonstrate higher correlation across varying benchmark sizes ($|X|$, denoted as $n_{\text{prompts}}$ in Figure 3).
-This empirical evidence further strengthens our argument for the effectiveness of the tournament approach.
-Spearman corr. ($\uparrow$) $|X|$ = 50 100 250 475 500
-anchored (4o) 0.895 0.935 0.963 0.966 0.964
-tournament (4o) 0.905 0.940 0.960 0.970 0.970
-anchored (4o-mini) 0.895 0.908 0.917 0.916 0.912
-tournament (4o-mini) 0.901 0.919 0.931 0.933 0.933
-Table 1: Spearman correlation ($\uparrow$) varying over size of the benchmark set ($|X|$) for each benchmarking approach.
-Values shown are median from 500 trials.
-See Figure 3 for detailed whisker plots.
-Varco Arena (labeled as tournament) consistently demonstrates higher correlation with the ground-truth leaderboard, regardless of deploying gpt-4o or gpt-4o-mini as the judge.
-Using the less expensive gpt-4o-mini as the judge did not compromise much in terms of correlation compared to its costlier counterpart.
-Notably, this tournament approach achieves comparable or superior correlation without the need for reference texts, while also reducing the number of required comparisons for benchmarking by $|X|$, further saving costs.
-The advantage of benchmarking by tournaments becomes more pronounced as the benchmark dataset size, $|X|$, increases.
-5.3 Incorporating a New LLM into an Existing Leaderboard
-While our main focus has been on ranking multiple LLMs simultaneously, it’s worth considering how our approach might apply to the common scenario of adding a single new model to an existing leaderboard.
-We explored a binary search-like placement method for this purpose (detailed in Appendix 2), but found it yielded less stable results than the reference output approach (Table 2).
-We attribute this instability to two factors: (1) The reduced number of comparisons (only $|X|$ comparisons per model, the minimum for a tournament run).
-(2) Incomplete coverage of the benchmark dataset’s prompts.
-However, the concept of a reference model as an anchor could still be valuable in this context.
-One could estimate the win rate against a anchor of one’s own choice for all models in the leaderboard, using this as a ranking score.
-For leaderboards built with Varco Arena, any missing matches could be imputed using expected relative win rates derived from Elo ratings or directly from existing match records.
-This approach offers a potential compromise between the stability of reference-based methods and the comparative advantages of our tournament approach.
-$\bar{|\Delta_{\text{rank}}|}$ ($\downarrow$) gt=1-6 7-13 14-19 (20) total avg.
-binary search (4o) 0.92 1.84 2.13 1.72
-anchored (4o) 1.98 1.55 1.57 1.39
-binary search (4o-mini) 1.27 1.82 1.21 1.5
-anchored (4o-mini) 1.00 1.43 1.43 1.37
-Table 2: Binary search vs.
-Anchored comparison for the insertion of a new LLM: Mean rank deviation ($\bar{|\Delta_{\text{rank}}|}$) from ground-truth leaderboard.
-Result of binary search placement and anchored comparison insert by gpt-4o[-mini] judge are provided with bootstrapped 95% confidence interval (500 trials, 1000 samples, $|X|$=500, Arena-Hard-Auto).
-6 Related Works
-6.1 Elo Modeling for LLM Benchmarking
-Utilizing Elo modeling of LLM capability based on human preference is a recently recognized means for benchmarking (Boubdir et al.
-2023).
-Chatbot Arena (Chiang et al.
-2024) is introduced as a benchmark platform that targets live benchmarking of LLMs toward open-ended generation tasks with the proper use of Elo ratings.
-Except that the platform requires massive human judge annotation, Chatbot Arena is considered a successful example of dynamic benchmarking for open-ended tasks as suggested in (Kiela et al.
-2021).
-RAGElo (Rackauckas, Câmara, and Zavrel 2024) deploy Elo model with LLM as a judge (Zheng et al.
-2023) to improve retrieval augmented generation (RAG) pipeline which indicates that modeling LLM capability with Elo rating is practically effective.
-6.2 Reference-free Evaluation
-Alleviating the need for reference text has long been considered an attractive goal in Natural Language Generation (NLG) evaluation.
-Recent advances in LLMs have shown promising results in their ability to assess open-ended responses (Jauhiainen and Guerra 2024).
-Also, findings from the WMT23 metrics task in neural machine translation (Freitag et al.
-2023) have demonstrated that reference-free quality estimation metrics outperform reference-based metrics in terms of correlation to human evaluation.
-Specifically, metrics such as XComet (Guerreiro et al.
-2023) and MetricX (Juraska et al.
-2023) have shown superior performance, particularly when dealing with scenarios involving poor-quality references.
-7 Conclusion
-We proposed Varco Arena, a reference-free LLM benchmarking approach for measuring generative capabilities of LLMs.
-Varco Arena deploys a tournament-style approach with direct pairwise comparisons, eliminating the need for reference outputs.
-Our simulation experiments demonstrate that this tournament approach achieves aligned LLM rankings better than current practice of reference use, under unbiased judging conditions.
-We also report the empirical effectiveness of Varco Arena using gpt-4o[-mini] as a judge, which reaffirms the efficacy of our approach in practice by showing improved correlation with Chatbot Arena leaderboard rankings.
-Through extensive experimentation and rigorous statistical analysis, we strongly suggest that our reference-free tournament approach, Varco Arena, provides a superior solution for LLM benchmarking, requiring fewer computational resources.
-Appendix A Appendix
-A.1 Machine Requirements for Experiments
-All the experiments we performed are conducted by querying API and post-processing those, which does not require GPU machine locally.
-Experiments could be run on personal desktops.
-The lowest specification of the machine we deployed had i5-8400 CPU, 16 GiB RAM.
-Figure 4: Elo ratings of the model with gpt-4o judge on the full set of Arena-Hard-Auto (Li et al.
-2024) prompts.
+Figure 4: Elo ratings of the model with gpt-4o judge on the full set of Arena-Hard-Auto Li et al.
+(2024) prompts.
 VARCO Arena result (bootstrapped median over 1000 samples of 500 trials) is in blue, plotted alongside the ratings from the ground truth leaderboard in red (Chatbot Arena, Hard prompts category).
 Error bars are 95% confidence intervals.
-A.2 Elo ratings from Varco Arena compared to Human Annotations
+A.3 Elo ratings from VARCO Arena compared to Human Annotations
 Figure 4 shows the Elo ratings computed out of VARCO Arena.
@@
 Matches are performed over Arena-Hard-Auto benchmark dataset (500 prompts).
-A.3 Binary search vs. Win rate over reference
+A.4 Binary search vs. Win rate over reference
 Algorithm 2 Binary Search for Enlisting new LLM to a leaderboard
@@
 35:end function
-Binary Search
+A.4.1 Binary Search
 We tried binary search placement of a newly added LLM to the leaderboard without reference text in Table 5.
@@
 Within the size of the benchmark prompts ($|X|$), binary search is incompatible with the current approach of using reference instead.
-Converting Elo Table back to Win Rate
+A.4.2 Comparing to the most performant Model so far: Converting Elo Table back to Win Rate
 Assuming we preserved a set of match results and model outputs from the last benchmarking, we could benefit from those to perform insertion.
@@
 Anchored comparison: Mean rank deviation ($|\Delta_{\text{rank}}|$) from ground-truth leaderboard.
-Result of binary search placement and anchored comparison insert by gpt-4o[-mini] judge are provided with bootstrapped 95% confidence interval (500 trials, 1000 samples, $|X|$=500, Arena-Hard-Auto (Li et al.
-2024)).
-A.4 Separability In terms of Confidence Interval
+Result of binary search placement and anchored comparison insert by gpt-4o[-mini] judge are provided with bootstrapped 95% confidence interval (500 trials, 1000 samples, $|X|$=500, Arena-Hard-Auto Li et al.
+(2024)).
+A.5 Separability In terms of Confidence Interval
 To see how well the two benchmarking approach (anchored comparison and tournament approach) separates LLMs in adjacent ranks, we provide scatter plot of Elo rating and win rate paired with error bar (95% confidence interval).
@@
 Figure 6: gpt-4o result of anchored comparison and tournament approach. 1000 bootstrapped median from 500 observations used for confidence interval estimation.
-A.5 Judge configuration
-Evaluation Prompt
+A.6 Judge configuration
+A.6.1 Evaluation Prompt
 We use the prompt from LLMBar.
-The prompt depicted in Figure 1.
-We added 4 questions for criteria of our own to Metrics.txt prompt of (Zeng et al.
-2024).
+The prompt depicted in Figure A.6.2.
+We added 4 questions for criteria of our own to Metrics.txt prompt of Zeng et al.
+(2024).
 You can refer to the original prompt in LLMBar github.
-Decoding Parameters
+A.6.2 Decoding Parameters
 We did not configure decoding parameters of judge LLMs (gpt-4o[-mini]), which its temperature defaults to 1.
@@
 To avoid position bias, we alternated the position of the responses from a certain model across the benchmark prompt.
-Listing 1: LLMBar prompt (Metrics.txt)
-1PROMPTS = [
-2 {
-3 role”: system”,
-4 content”: ”””You are a helpful assistant in evaluating the quality of the outputs for a given instruction. Your goal is to select the best output for the given instruction.”””,
-5 },
-6 {
-7 role”: user”,
-8 content”: ”””Select the Output (a) or Output (b) that is better for the given instruction. The two outputs are generated by two different AI chatbots respectively.
-9
-10Here are some rules of the evaluation:
-11(1) You should prioritize evaluating whether the output honestly/precisely/closely executes the instruction, then consider its helpfulness, accuracy, level of detail, harmlessness, etc.
-12(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
-13(3) You should avoid any potential bias and your judgment should be as objective as possible.
+PROMPTS = [ # metrics.txt from LLMBar
+"role": "system", "content": "You are a helpful assistant in evaluating the quality of the outputs for a given instruction. Your goal is to select the best output for the given instruction.", ,
+"role": "user", "content": """Select the Output (a) or Output (b) that is better for the given instruction. The two outputs are generated by two different AI chatbots respectively.
+Here are some rules of the evaluation:
+(1) You should prioritize evaluating whether the output honestly/precisely/closely executes the instruction, then consider its helpfulness, accuracy, level of detail, harmlessness, etc.
+(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
+(3) You should avoid any potential bias and your judgment should be as objective as possible.
 For example, the order in which the outputs were presented should NOT affect your judgment, as Output (a) and Output (b) are **equally likely** to be the better.
-14
-15Do NOT provide any explanation for your choice.
-16Do NOT say both / neither are good.
-17You should answer using ONLY Output (a)” or Output (b)”. Do NOT output any other words.
-18
-19# Instruction:
-20{instruction}
-21
-22# Output (a):
-23{response_a}
-24
-25# Output (b):
-26{response_b}
-27
-28# Questions about Outputs:
-29Here are at most three questions about the outputs, which are presented from most important to least important. You can do the evaluation based on thinking about all the questions.
-30* Does the output well satisfy the intent of the user request?
-31* If applicable, is the output well-grounded in the given context information?
-32* Does the output itself satisfy the requirements of good writing in terms of:
-33 1) Coherence
-34 2) Logicality
-35 3) Plausibility
-36 4) Interestingness
-37
-38
-39# Which is better, Output (a) or Output (b)? Your response should be either Output (a)” or Output (b)”:”””,
-40 },
-41]
-Appendix B Reproducibility Checklist
-  • Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes)
-  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes)
-  • Provides well marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper (yes)
-  • Does this paper make theoretical contributions? (yes)
-  • All assumptions and restrictions are stated clearly and formally. (yes)
-  • All novel claims are stated formally (e.g., in theorem statements). (partial)
-  • Proofs of all novel claims are included. (partial)
-  • Proof sketches or intuitions are given for complex and/or novel results. (yes)
-  • Appropriate citations to theoretical tools used are given. (yes)
-  • All theoretical claims are demonstrated empirically to hold. (yes)
-  • All experimental code used to eliminate or disprove claims is included. (no)
-  • Does this paper rely on one or more datasets? (yes)
-  • A motivation is given for why the experiments are conducted on the selected datasets (yes)
-  • All novel datasets introduced in this paper are included in a data appendix. (NA)
-  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes)
-  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. (yes)
-  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. (yes)
-  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. (NA)
-  • Does this paper include computational experiments? (yes)
-  • Any code required for pre-processing data is included in the appendix. (no)
-  • All source code required for conducting and analyzing the experiments is included in a code appendix. (no)
-  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes.
-(yes)
-  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (partial)
-  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. (yes)
-  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks.
-(partial)
-  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. (yes)
-  • This paper states the number of algorithm runs used to compute each reported result. (yes)
-  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. (yes)
-  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). (yes)
-  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes)
-  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes)
+Do NOT provide any explanation for your choice.
+Do NOT say both / neither are good.
+You should answer using ONLY "Output (a)" or "Output (b)". Do NOT output any other words.
+# Instruction:
+instruction
+# Output (a):
+response_a
+# Output (b):
+response_b
+# Questions about Outputs:
+Here are at most three questions about the outputs, which are presented from most important to least important. You can do the evaluation based on thinking about all the questions.
+- Does the output well satisfy the intent of the user request?
+- If applicable, is the output well-grounded in the given context information?
+- Does the output itself satisfy the requirements of good writing in terms of:
+1) Coherence
+2) Logicality
+3) Plausibility
+4) Interestingness
+# Which is better, Output (a) or Output (b)? Your response should be either "Output (a)" or "Output (b)":""",